### v1

# =COLI-manual3.qry Cross-lingual Cross-temporal Summarization: Dataset, Models, Evaluation

Abstract
While summarization has been extensively researched in natural language processing (NLP), cross-lingual cross-temporal summarization (CLCTS) is a largely unexplored area that has the potential to improve cross-cultural accessibility, information sharing, and understanding. This paper comprehensively addresses the CLCTS task, including dataset creation, modeling, and evaluation. We build the first CLCTS corpus, leveraging historical fictive texts and Wikipedia summaries in English and German, and examine the effectiveness of popular transformer end-to-end models with different intermediate task finetuning tasks. Additionally, we explore the potential of ChatGPT for CLCTS as a summarizer and an evaluator. Overall, we report evaluations from humans, ChatGPT, and several recent automatic evaluation metrics where we find our intermediate task finetuned end-to-end models generate bad to moderate quality summaries; ChatGPT as a summarizer (without any finetuning) provides moderate to good quality outputs and as an evaluator correlates moderately with human evaluations though it is prone to giving lower scores. ChatGPT also seems to be very adept at normalizing historical text. We finally test ChatGPT in a scenario with adversarially attacked and unseen source documents and find that ChatGPT is better at omission and entity swap than negating against its prior knowledge.
1 Introduction
Summarization is a key task in natural language processing (NLP), especially in an age of information overload. The classical approach for summarization is to summarize documents (e.g., news reports) in the same language Zhang et al. (2020a); Liu et al. (2022); Ravaut, Joty, and Chen (2022) or, less prominently, in different languages than the original source (cross-lingual summarization; CLS) Ladhak et al. (2020); Cao, Liu, and Wan (2020); Bai, Gao, and Huang (2021); Liang et al. (2022). Tasks such as the summarization of historical texts to modern languages (cross-temporal summarization; CTS) are scarcely explored, in contrast. Such summarizers are beneficial not only for historical researchers and students but also for laypeople with an interest in historical information sources.
In this work, we go one step further and explore cross-lingual cross-temporal summarization (CLCTS), in which the goal is to summarize a historical document in a different modern language. See Table 1 for an example. Such cross-lingual and cross-temporal summarizers would potentially tremendously facilitate cross-cultural accessability, information sharing, and understanding.
Title |
An Inhabitant of Carcosa |
Author |
Ambrose Bierce |
Year |
1886 |
Source |
The Collected Works of Ambrose Bierce, Volume 3 |
Text |
“For there be divers sorts of death—some wherein the body remaineth; and in some it vanisheth quite away with the spirit … In one kind of death the spirit also dieth, and this it hath been known to do while yet the body was in vigor for many years. […]” |
Summary (German) |
“Ein Mann aus der Stadt Carcosa, der über die Worte des Philosophen Hali über die Natur des Todes nachdenkt, wandert durch eine ihm unbekannte Wildnis. […]” |
CLCTS is difficult and challenging for the following reasons: (i) languages change over time along multiple dimensions, including syntax Juzek, Krielke, and Teich (2020); Lei and Wen (2020), semantics Hamilton, Leskovec, and Jurafsky (2016); Giulianelli, Del Tredici, and Fernández (2020), lexical choice and morphology Gibson et al. (2019); Joseph (2017). (ii) Historical documents are often longer, not only because of tendencies of simplification/text length reduction over time Lewis (1894); Sherman (1893); Rudnicka (2018), but also because of the genre of historical text, which predominantly includes literary work Wang and Liu (2017); Zhu, Liu, and Pang (2022). (iii) The process can be thought of as consisting of several subtasks, previously considered independent in NLP research: machine translation (MT), summarization and historical (spelling) normalization Bollmann and Søgaard (2016); Eger, vor der Brück, and Mehler (2016).
In this work, we consider the CLCTS problem comprehensively, including dataset creation, modeling, and evaluation.111Code+data: https://github.com/zhangr2021/CLCTS We start with building the first CLCTS dataset. Our corpus contains stories and plays (as prime artefacts of historical cultures), which requires the models to handle long documents; we assemble modern summaries by querying Wikipedia.
In terms of modeling, we consider three approaches: (1) extractive summarizers Gu, Ash, and Hahnloser (2022) whose advantage is that they can in principle deal with unlimited input document lengths but which are limited in that they can only copy out sentences from the original source (which is a problem, e.g., when the source is first-person narrative but the summary should be third-person perspective). (2) Abstractive summarizers can avoid the issue above but typically have severe limitations in input lengths. Even models exclusively designed for long document summarization by utilizing efficient variants of attention mechanism such as Longformer Encoder-Decoder (LED) Beltagy, Peters, and Cohan (2020) and Bigbird Zaheer et al. (2020) limit the max input tokens to 16,384 and 4,096, respectively. (3) GPT-3 Brown et al. (2020) and its extensions such as ChatGPT are also strong candidates for our task Goyal, Li, and Durrett (2022) as they have so spectacularly impacted research in a multitude of scenarios Leiter et al. (2023), including summarization Yang et al. (2023); Wang et al. (2023); Zhang, Liu, and Zhang (2023). Consequently, we also include ChatGPT in our analysis.
In terms of evaluation, we consider three evaluation methods: (1) Automatic evaluation, where we report multiple evaluation metrics additional to the commonly used variants of ROUGE Lin (2004) in summarization tasks Liang et al. (2022); Bai, Gao, and Huang (2021); Cao, Liu, and Wan (2020); Peng et al. (2021) since the performance of individual metrics may vary across datasets, challenging their reliability Bhandari et al. (2020); Fabbri et al. (2021). (2) Human evaluation, which reflects the actual quality of summaries according to human judgements and functions as a source of reliability measurement for automatic evaluation metrics. (3) ChatGPT evaluation, where we examine the potential of ChatGPT as an alternative to human annotators given the same instruction considering the high cost of human evaluation Gao et al. (2023) and issues of reproducibility Belz et al. (2021); Belz, Thomson, and Reiter (2023); Chen, Belouadi, and Eger (2022).
Our contributions are:
To our best knowledge, we build the first CLCTS corpus, leveraging historical fictive texts and Wikipedia summaries.
We examine the effectiveness of popular transformer end-to-end (e2e) models with different tasks for intermediate task finetuning Chang and Lu (2021).
We explore how ChatGPT performs for CLCTS.
We provide both human, ChatGPT and automatic evaluation for the task, showing that our intermediate task finetuned e2e models generate bad to moderate quality summaries; ChatGPT as a summarizer provides moderate to good quality outputs (without any finetuning); ChatGPT as an evaluator correlates with human evaluations to a moderate level though it is prone to lower scores.
2 Related work
[…]
Since we focus on various summarization instead of translation and long documents instead of paragraphs,
the translation dataset is not within the scope of our current study.
Cross-lingual summarization
[…]
Long document summarization is another area of our concern.
[…]
attention (input limits 4096 and 16384) and PEGASUS with BigBird attention (4096).333The very recently proposed retrieval-based approach unlimiformer Bertsch et al. (2023) extends the input limitations of the aforementioned pretrained encoder-decoder transformers to practically unlimited input sequences by offloading the cross-attention computation.
We build our model with mBART
Tang et al. (2020) pretrained for multilingual MT tasks and combine with LED attention to enable efficient processing for long sequences.
Additionally, we build another retrieve-then-summarize pipeline
where we retrieve sentences with an extractive summarizer Gu, Ash, and Hahnloser (2022) and then summarize cross-lingually using ChatGPT.
ChatGPT for summarization Despite ChatGPT’s young age, there is already a wealth of research that explores it for various NLP tasks, including summarization Yang et al. (2023) and summarization evaluation Chiang and yi Lee (2023); Gao et al. (2023). Disagreement exists concerning ChatGPT’s performance.
Most studies focus on MLS tasks.
[…]
They also claim that ChatGPT can perform better with pipeline-like prompt lines
in an interactive manner. It is also worth noticing that all works mentioned above utilize small samples of 50 to 100 random samples (out of 3000 to 60k observations) from each dataset in their experiments.
Summarization evaluation metrics Depending on the availability of reference summaries, one can utilize reference-based metrics (focusing on overlap measure Lin (2004); Popović (2015), embedding similarity Zhang et al. (2020b); Zhao et al. (2019), discourse coherence Zhao, Strube, and Eger (2023), natural language inference (NLI) Chen and Eger (2023) and text generation task such as question answering (QA) Yuan, Neubig, and Liu (2021); Wang, Cho, and Lewis (2020)) or reference-free metrics Gao, Zhao, and Eger (2020); Chen, Li, and King (2021); Liu, Jia, and Zhu (2022); Laban et al. (2022); Belouadi and Eger (2023).
[…]
To address this concern, we leverage multiple automatic evaluation metrics, especially more recent ones e.g., MENLI and DiscoScore, as well as human evaluation. Since we focus on long documents,
we do not explore reference-free metrics, due to the vast mismatch in length between source texts and summaries.
[…]
3 Dataset
In this section, we introduce our CLCTS corpus and additional sources we utilize for our experiments. We build our own corpus which covers English and German in both directions. We utilize multiple monolingual and cross-lingual summarization datasets as additional resources during our intermediate task finetuning process.
3.1 CLCTS Corpus
Manually creating summaries is very time-consuming and requires expert knowledge Peng et al. (2021). Thus, we use fairy tales, short stories and plays where summaries are available in Wikipedia articles. In each language direction, English and German, the dataset contains summary pairs from historical documents in one language to modern summaries in the other language.
Dataset collection We collect historical documents mainly from three different sources. The links to the sources are collected in Appendix 9.1.
Deutsches Textarchiv (DTA; German text archive): This is a basic stock of German-language texts focusing on the early 16th to early 20th centuries.
Wikisource: This is a multilingual online digital library of free-textual content containing historical books in text format and original format stored as images. Some characters are changed from the original historical format to modern digital formats. Proofreading is performed for the texts by users on Wikisource following the proofreading guide.
Project Gutenberg: This is a cultural digital archive with over 60,000 digitized books, including historical texts. It focuses on older works with expired U.S. copyrights.
Our workflow for building the corpus consists of two stages: (1) meta-information collection & historical text extraction and (2) summary collection & translation. In stage (1), we collect meta-information of historical documents using the “Beautifulsoup” library. We extract information such as titles, web links, authors, year, and documents together with information from Wikipedia lists of German fairy tales and English short stories. For documents presented in several sources, the oldest version or the version with historical spelling is selected. For stage (2), we utilize the collected information to match the historical documents with their corresponding summaries from Wikipedia. The historical spelling of titles is converted to modern spelling to match the titles in Wikipedia. Figure 1 illustrates the process of data collection. If no original human correspondent summaries are found in the other language, we translate the summaries into the target language using DeepL (we translate 144 out of 328 German summaries to English and 108 out of 289 English summaries to German).555DeepL API: https://www.deepl.com/pro-api An example of our collected dataset is shown in Table 1.
In addition to the summarization datasets, we create a historical German-English translation dataset with 201 translation pairs. This dataset is later utilized in intermediate task finetuning. We retrieve the German documents from the DTA corpus and the historical English from Wikisource. The documents are translated by Margaret Raine Hunt (1831–1912). An example of the translation dataset is in Table 13 (appendix).
We use the abbreviations hDe and hEn for historical German and historical English texts to distinguish them from modern texts De or En. Thus, we refer to the historical German to modern English summarization dataset as “hDe-En” and the historical English to modern German summarization dataset as “hEn-De”. Similarly, monolingual CTS datasets are called “hDe-De” or “hEn-En”. Names of other modern datasets follow the same rule.
3.2 External datasets
To enrich our finetuning data resources, we utilize external summarization datasets: (i) HISTSUMM Peng et al. (2021) where we also translate the summaries into English using DeepL to enable cross-lingual finetuning. (ii) We further use Wikilingua for MLS and CLS tasks Ladhak et al. (2020). Wikilingua is a multilingual summarization dataset which contains 18 languages from WikiHow (an online wiki-style publication featuring how-to articles), including English and German. The English dataset contains 141,457 document-summary pairs. For finetuning our models, we randomly select 60% of the pairs (85,874) and we use all of the German dataset (58,341). Wikilingua German entries contain the URL and the section name of the English articles. Thus, cross-lingual summary pairs can be retrieved using this information and we match 20,103 cross-lingual summary pairs in both directions. (iii) We also utilize the CNN/Daily Mail dataset. This is a monolingual English summarization dataset containing 311,971 document-summary pairs, including news articles written by CNN and Daily Mail journalists, and summaries from the highlight of the article written by the article author. We experiment with a subset of 86,133 text-summary pairs. (iv) We include MLSUM (Scialom et al., 2020), which is a multilingual summarization dataset containing over 1.5 million document-summary pairs collected from online newspapers such as Süddeutsche Zeitung (German). We use a subset of 66,226 pairs from the German dataset.
3.3 Dataset Statistics
Dataset | Size | Mean Length | Compression | Genre | Task | |||
Doc. | Summ. | |||||||
CLCTS Corpus | ||||||||
CLCTS hEn-De | 289 | 9,643.2 | 413.9 | 16.0 | 19.2 | 23.3 | fiction | CLS, CTS |
CLCTS hDe-En | 328 | 1,416.8 | 324.1 | 25.4 | 18.6 | 4.4 | fiction | CLS, CTS |
CLCTS hEn-En | 289 | 9,643.2 | 423.6 | 16.0 | 20.2 | 22.8 | fiction | MLS, CTS |
CLCTS hDe-De | 328 | 1,416.8 | 217.3 | 25.4 | 16.6 | 6.5 | fiction | MLS, CTS |
External Corpus | ||||||||
HISTSUMM hDe-De | 100 | 268.1 | 18.1 | 36.3 | 12.4 | 14.8 | news | MLS, CTS |
Wikilingua En-En | 85,874 | 425.8 | 39.6 | 16.5 | 6.5 | 10.8 | how-to-guide | MLS |
Wikilingua De-De | 58,341 | 429.8 | 42.2 | 15.8 | 6.9 | 10.2 | how-to-guide | MLS |
Wikilingua De-En | 20,103 | 438.2 | 38.8 | 15.8 | 6.5 | 11.3 | how-to-guide | CLS |
Wikilingua En-De | 20,103 | 451.7 | 42.7 | 16.5 | 6.9 | 10.6 | how-to-guide | CLS |
CNN/Daily Mail En-En | 86,133 | 786.7 | 55.1 | 20.5 | 14.5 | 14.3 | news | MLS |
MLSUM De-De | 66,226 | 570.3 | 30.4 | 17.8 | 12.7 | 18.8 | news | MLS |
In Table 2, we report statistics of our corpus and other existing summarization datasets. We observe three trends comparing all datasets:
Much fewer instances in Historical datasets. The CLCTS hEn-De dataset consists of 289 historical short stories and plays, and the CLCTS hDe-En dataset consists of 328 historical fairy tales. The size of HISTSUMM dataset is even smaller, with only 100 historical news articles. In contrast, modern summarization datasets contain several thousand summarization pairs, which is more than 100 times the size of historical datasets. The largest dataset CNN/Daily Mail contains over 300k summarization pairs.
Longer documents and summaries for CLCTS corpus. The CLCTS hEn-De dataset has more than 9643 tokens per document and 413 tokens per summary on average and CLCTS hDe-En has 1416 per document and 324 per summary. For external datasets, the average length of documents is below 800 tokens per document and below 60 tokens per summary. This is because the CLCTS corpus includes historical documents from the genre fiction, which are naturally longer than news and WikiHow (how-to guide) due to genre differences (Rudnicka, 2018) and higher verbosity of more historical languages.
Larger information loss during summarization for news genre and CLCTS hEn-De measured by compression ratio. News documents have a higher compression ratio (the mean length of document divided by the mean length of summary) compared to documents from other genres, except for CLCTS hEn-De, which has the highest compression ratio among all datasets. The CLCTS hEn-De dataset contains short stories and plays which are much longer than fairytales in CLCTS hDe-En and documents from all other genres as discussed above. For the CLCTS corpus, the compression ratio differs substantially because all summaries of the CLCTS corpus are collected from Wikipedia with the average length of summaries ranging from 200 to 400 tokens regardless of the original document length.
To obtain more information on document-summary pairs, we quantify the divergence of sentence semantics between document and summary by computing the mean sentence-level cosine similarity for all datasets (Vyas, Niu, and Carpuat, 2018). We obtain the sentence embedding via multilingual SBERT (Reimers and Gurevych, 2020). We first compute the average similarity of each source to each summary sentence to obtain document level similarity and then average over all documents to obtain dataset level similarity score. The results are shown in Table 3. It reveals that cross-lingual sentence semantics between document and summary is more diverged than that of monolingual datasets and among all, both CLCTS hDe-En and hEn-De (which are our target tasks) have lower embedding similarity compared to other datasets. This again indicates the difficulty of cross-lingual and cross-temporal summarization.
Monolingual | Cross-lingual | ||
CLCTS hDe-De | 0.38 | CLCTS hDe-En | 0.32 |
CLCTS hEn-En | 0.27 | CLCTS hEn-De | 0.25 |
HISTSUMM hDe-De | 0.39 | - | - |
CNN/Daily Mail En-En | 0.32 | - | - |
MLSUM De-De | 0.37 | - | - |
Wikilingua De-De | 0.46 | Wikilingua De-En | 0.41 |
Wikilingua En-En | 0.42 | Wikilingua En-De | 0.42 |
We include an analysis of mean sentence length in the next section where we focus on a detailed analysis of historical documents and their comparison to modern documents.
3.4 Historical Language Divergence
Languages evolve due to social, cultural and linguistic pressure and this poses challenges to the summarization of historical texts.
Publication time of historical documents
The distribution of publication year is shown in Figure 2. HISTSUMM documents (in red) distribute evenly between the time period of 1650–1800. CLCTS hEn-De (in gray) contains more documents from the late 1800s and CLCTS hDe-En (in green) from the early and mid 1800s. The documents from CLCTS hEn-De which date back to the 1600s are plays from Shakespeare. For both CLCTS datasets, the majority of historical documents come from the 1800-1900s and comparatively HISTSUMM has overall older documents than our CLCTS corpus.
Divergence in syntax Length of Sentences The analysis in Table 2 (column 5) reveals that historical German documents in HISTSUMM and CLCTS corpus contain longer sentences compared to other datasets, where HISTSUMM hDe-De has a mean sentence length of 36.3 tokens and CLCTS hDe-En/hDe-De dataset has 25.4 tokens per sentence while modern German documents from Wikilingua and MLSUM have maximum 20 tokens per sentence on average. This coincides with the finding of decreasing sentence length over time (Rudnicka, 2018), considering that HISTSUMM documents are the oldest. Slightly surprisingly, this is not true for the English documents, where the mean sentence length of historical English documents (CLCTS hEn-De/hEn-En) is almost the same as Wikilingua and is even shorter than documents from news genre CNN/Daily Mail. This hints at a genre effect, besides the temporal effect: fiction, especially plays, may have shorter sentences than news and other non-fiction genres due to the higher occurrence of dialogues and two-word sentences (Rudnicka, 2018). In our CLCTS corpus, unlike CLCTS hDe-En/hDe-De containing only fairytales, historical English documents contain plays. When we exclude plays from Shakespeare, the mean sentence length of the CLCTS hEn-De document increases by 4 tokens per sentence to 20.3, which is closer to the value of CNN/Daily Mail. The factual summaries from CLCTS hEn-De/hEn-En in column 6 also contain longer sentences than their fictional documents for similar reasons.
Dependency Distances To further quantify syntactic changes, we analyze the dependency parsing trees for sentences from documents in all datasets and compute the mean dependency distance (MDD) over the entire corpus Liu (2008); Liu, Xu, and Liang (2017).
This measure can reflect the syntactic complexity of sentences. We use stanza (Qi et al., 2020) for parsing the sentences. We plot the
MDD
in Figure 3 where we observe a downward trend overall. The first four documents from the left are in German and the last three documents are in English (we use dataset plus language e.g., CLCTS hDe to represent the documents here). The downward trending line reveals an overall higher MDD of German sentences than that of English sentences. If we take a closer look by language, we observe: (i) For German (the first four points from the left), historical sentences (in red star) have a higher value for MDD than that of the modern German text, which suggests more syntactic complexity of historical documents. We especially notice a higher MDD from CLCTS hDe documents (fiction) than that from HISTSUMM hDe (news), which coincides with the finding of longer dependencies in the imaginative genre than those in the informative genre from Wang and Liu (2017). The modern sentences from MLSUM are also more complex syntactically compared to Wikilingua. (ii) For English (top-3 right items), this is not entirely true. News documents from CNN/Daily Mail have again a longer MDD than that of Wikilingua. However, they also have a longer MDD than CLCTS hEn.
This may be because of the combined effect of (a) shorter sentences from CLCTS hEn (also from Wikilingua En) as discussed above, (b) more conversational texts with shorter dependencies from CLCTS hEn documents (plays and short stories) Wang and Liu (2017) and (c) CLCTS hEn documents mostly from the late 1800s which share more similarities to modern English.
Divergence in lexicon The vocabulary used in historical and modern language varies greatly due to spelling, morphological and semantic changes. For example, in our historical German dataset, the letter “ß” often represents “ss” in modern texts (e.g., “daß” vs. “dass”). Omitting “h”, “t” or “e” is another type of spelling change such as “Thüre” (door, now “Türe” or “Tür”), “thun” (to do, now tun), “Todt” (death, now “Tod”) and “gieng” (passive tense of to go, now “ging”).
Such changes in German are a result of two spelling reformations, one in 1901 and one in 1996.
In English, we can also observe changes in lexicon such as the word “you”
written as “thou” and the morphological change of the second person “will” as “wilt”. To quantify these divergences, we compute the Jaccard similarity (Niwattanakul et al., 2013) of lexicons between documents and their summaries in Table 4. Historical German has lowest lexical overlap with modern German, with a Jaccard score of 0.1 for HISTSUMM and 0.19 for CLCTS hDe-De.
The pattern is similar for English.
Dataset | Jaccard |
CLCTS hEn-En | 0.234 |
CNN/Daily Mail En-En | 0.377 |
Wikilingua En-En | 0.342 |
CLCTS hDe-De | 0.186 |
HISTSUMM hDe-De | 0.101 |
MLSUM De-De | 0.335 |
Wikilingua De-De | 0.314 |
Semantic shift is another source of language change (Eger and Mehler, 2016; Hamilton, Leskovec, and Jurafsky, 2016). For example, the word “mistress” in Shakespeare’s work has a similar meaning as “sweetheart”. The “car” mentioned in Shakespeare’s work (e.g., “Phoebus’ car”) refers to a special kind of chariot. Another example is the adjective “gay” which is used in the sense of “glad and cheerful”, e.g., “The great city in which he dwelt was very gay, for every day strangers visited the town” (The Emperor’s New Clothes by Christian Andersen in 1888) . Similarly, the meaning of the word “knave” degenerates from “servant” in the story of brother Grimm to the modern meaning “a dishonest man” over time.
4 Experimental setup
Our experiments consist of three methods: pipeline methods based on extractive summarizers, e2e transformer-based methods with intermediate task finetuning and zero-shot ChatGPT with different prompt strategies. For finetuned methods (extractive and abstractive), we use 5-fold cross-validation to obtain reliable results for all CLCTS datasets. For ChatGPT, we query summaries for all CLCTS datsets as well to make sure the results are comparable.
4.1 Pipeline based on extractive summarizer
For extractive summarization, we first use an extract-then-translate pipeline. We choose MemSum (Gu, Ash, and Hahnloser, 2022) as our model, which utilizes multi-step episodic Markov decision processes to extract important sentences from a given text. We use 5-fold cross-validation to train and test MemSum with our monolingual dataset. We limit the maximal extracted sentences per text to 25 according to the average length of summaries in Section 3. During inference, we sort the sentences according to the original text’s order, then translate the extracted sentences to the target language using a multilingual machine translation model developed by Facebook (M2M) (Fan et al., 2021). We additionally experiment with historical text normalization before translation. We use the normalization method Norma666https://github.com/comphist/norma (Bollmann, 2012). We choose the combined version which leverages Lookup, Rule-based and Distance-based normalization which is claimed to perform well for both English and German (Bollmann, 2019).
To tackle input limitations of ChatGPT (e.g., for GPT-3.5-turbo a total of 4096 tokens are shared between prompt and completion777https://help.openai.com/en/articles/4936856-what-are-tokens-and-how-to-count-them), we build the third retrieve-then-summarize pipeline, where we use MemSum as a key sentence retriever and input the retrieved sentences to ChatGPT and request summarization in the target language. For this experiment, we limit the maximal extracted sentences per document to 100 according to the input limits of ChatGPT.
4.2 Transformer-based e2e models
We modify pretrained models together with intermediate task
finetuning to build e2e models. We propose a modified multilingual mBART50 model pretrained for machine translation Tang et al. (2020) with an efficient attention mechanism taken from LED (Beltagy, Peters, and Cohan, 2020) as model architecture. This architecture is suitable for multilingual long sequences as it benefits from the cross-linguality of mBART and the efficiency of processing long documents from LED. We refer to this model as mLED.
Baselines
We first provide baselines of mLED finetuned only within the CLCTS corpus: i)
for baseline 1, we finetune the model only with the target task; e.g., for the model in direction hDe-En, we only finetune mLED with our training set of the same direction. We refer to this model as Base.
ii)
For baseline 2 (baseline 1 + historical machine translation), we utilize our historical translation dataset. We first finetune with the translation task and then with the target task, which we call Historical MT.
iii) Finally, for baseline 3, we finetune with datasets from both directions. We add prefixes to indicate the directions. We refer to this model as .
Intermediate task finetuning
Apart from the final finetuning step with our target CLCTS task, we conduct intermediate task finetuning using external sources including MLS, CLS, and CTS. Table 5 depicts the detailed design. We use 5-fold cross-validation to obtain reliable results. Due to resource limitations, our mLED models have a maximum of 4,096 input tokens and the number of tokens for the generated summary ranges from a minimum length of 100 tokens to a maximum length of 512 tokens.
Model | MLS | CLS | CTS |
CNN/Daily Mail En-En (trg) | |||
MLSUM De-De (src) | |||
CNN/Daily Mail En-En (trg) | |||
MLS-CLS | Wikilingua De-De & En-En | Wikilingua De-En | |
MLS-CLS-CTS | Wikilingua De-De & En-En | Wikilingua De-En | HISTSUMM hDe-En |
(a) Direction hDe-En | |||
Model | MLS | CLS | CTS |
MLSUM De-De (trg) | |||
Wikilingua En-En (src) | |||
Wikilingua De-De (trg) | |||
MLS-CLS | Wikilingua En-En & De-De | Wikilingua En-De | |
MLS-CLS-CTS | Wikilingua En-En & De-De | Wikilingua En-De | HISTSUMM hDe-De |
(b) Direction hEn-De |
4.3 CLCTS with ChatGPT
We test the summarization ability of ChatGPT with our CLCTS corpus.
Prompt Strategies
We utilize GPT-3.5-turbo
via the OpenAI API to query the summary. We set the temperature to 0.7 to allow for some randomness. We adopt (i) a pipeline translate-then-summarize prompt as well as a (ii) e2e prompt. We set prompt lines to the same language as the target output since ChatGPT works better in this setting according to our initial experiments.888During our first round of data collection all the prompt lines were in English. We noticed that ChatGPT tends to output summaries in the wrong target language, especially when the summary is expected to be in German.
We truncate the texts that exceed ChatGPT input limits.999According to our experiment with the GPT2 tokenizer, we find that on average our German texts have 1.73 tokens per word (which is higher than the suggested 1.3 tokens per word ratio from OpenAI) and English 1.07 tokens per word. Therefore, we limit the German text to 2048 words and English to 3000 words. Since the 4096 tokens limits from GPT-3.5-turbo are shared between both prompts and completions (outputs), we need to leave some tokens for the output summary.
We also utilize other information such as title and author to test if ChatGPT has explicit “prior” knowledge of the input document and its corresponding summary. We list all our prompt lines in Table 6.
Experiment | Src-Trg | Prompt |
e2e | hDe-En |
Please summarize the following text in English : [text] |
e2e | hEn-De |
Bitte fasse den folgenden Text auf Deutsch zusammen: [Text] |
e2e (Title) | De-En |
Please give me the summary of the story [Title] written by [Author]. |
e2e (Title) | En-De |
Bitte gebe mir die Zusammenfassung der Geschichte [Titel] von [Autor]. |
pipeline | hDe-En |
Please first translate the following text into English and summarize the translated text: [Text] |
pipeline | hEn-De |
Bitte übersetze zuerst den folgenden Text auf Deutsch und fasse den übersetzten Text zusammen: [Text]. |
4.4 Evaluation
To identify the best model, we utilize multiple recently popular reference-based metrics. We also conduct a multi-phase
human evaluation to evaluate the model and test the effectiveness of the resulting metrics. During the multi-phase evaluation, we give feedback to annotators after each phase for them to better align with the instructions.
Automatic Evalution
We take the ROUGE-variant Lin (2004) which measures the lexical overlap by matching the texts through computing the unigram overlap (ROUGE-1), and ROUGE-L, which leverages the longest common subsequence. BERTScore Zhang et al. (2020b) and MoverScore Zhao et al. (2019) focus on semantic overlap, where BERTScore calculates the token similarity and Moverscore utlizes Word Mover’s distance. We also consider generation-based BARTScore Yuan, Neubig, and Liu (2021), NLI-based MENLI Chen and Eger (2023) and discourse-based DiscoScore Zhao, Strube, and Eger (2023). The configurations used for each metric are presented in Table 14 (appendix).
Human Evaluation
We use six ‘expert’ annotators to evaluate all models.
Three of the annotators are female and three are male.
All annotators have high competence in English as they are affiliated with the university: they are Master students (3), postgraduates (1), PhD students (1) and NLP faculty members (1).
Three of them are native German speakers. Thus, the generated English summaries are evaluated by all six annotators and the generated German summaries are annotated by three native speakers.
We utilize the evaluation strategy by Fabbri et al. (2021) where the annotators score the summary from four perspectives: coherence, consistency, fluency, and relevance. The outputs are rated with a score ranging from 1 (worst) to 5 (best) for each dimension and we allow 0.5 increments. After phase 1, we give feedback to the annotators by providing a comparison of their scores with the average scores. The scores that differ substantially are
highlighted and we encourage them to review the differences.
To speed up the process, we did not evaluate the same amount of documents for all the models.
Nonetheless, we make sure that each model is evaluated for at least 4 identical source documents.
ChatGPT evaluation
Since human annotation is costly, we follow related work Gao et al. (2023); Chiang and yi Lee (2023) and test ChatGPT as an evaluator for our CLCTS task. We differ in the following aspects: 1) our CLCTS task is harder to evaluate due to the long length of both reference and generated summary. 2) Our CLCTS task involves multilingual evaluation for both English and German. We give the same instructions to ChatGPT as to the human annotators.
For reproducibility purposes, we set the temperature to 0 to reduce the randomness.
We collect annotations of generated summaries from all models from the documents annotated by humans, resulting in 501 ChatGPT annotations for hDe-En and 384 ChatGPT annotations for hEn-De.
5 Results
5.1 Automatic evaluation metrics
Model | ROUGE-1/L | BERTScore-P/R/F1 | BARTScore | MoverScore | MENLI-/// | DiscoScore |
Supervised Extractive | ||||||
MemSum | ||||||
- | 0.318/0.174 | 0.520/0.572/0.544 | -3.315 | 0.552 | -0.234/0.406/0.460/0.471 | 0.551 |
- | 0.314/0.171 | 0.513/0.566/0.537 | -3.363 | 0.551 | -0.250/0.396/0.448/0.459 | 0.536 |
- | 0.382/0.234 | 0.640/0.596/0.615 | -3.419 | 0.571 | -0.207/0.446/0.565/0.589 | 0.883 |
Supervised Abstractive - Baseline CLCTS | ||||||
mLED | ||||||
-Base | 0.391/0.201 | 0.537/0.561/0.547 | -3.524 | 0.568 | -0.206/0.418/0.469/0.479 | 1.838 |
-Historical MT | 0.393/0.200 | 0.539/0.566/0.551 | -3.492 | 0.569 | -0.216/0.416/0.473/0.485 | 1.725 |
0.388/0.198 | 0.537/0.561/0.547 | -3.534 | 0.567 | -0.231/0.408/0.465/0.477 | 1.142 | |
Supervised Abstractive - intermediate finetuning | ||||||
mLED | ||||||
0.389/0.198 | 0.540/0.556/0.547 | -3.567 | 0.568 | -0.255/0.398/0.461/0.473 | 1.426 | |
0.386/0.201 | 0.536/0.547/0.54 | -3.593 | 0.568 | -0.252/0.397/0.453/0.464 | 1.595 | |
-MLS-CLS | 0.380/0.198 | 0.530/0.544/0.536 | -3.615 | 0.567 | -0.241/0.400/0.448/0.458 | 1.685 |
-MLS-CLS-CTS | 0.386/0.202 | 0.532/0.548/0.539 | -3.590 | 0.568 | -0.246/0.399/0.451/0.462 | 1.635 |
Zero-shot Abstractive | ||||||
ChatGPT as summarizer | ||||||
-title prompt | 0.304/0.164 | 0.527/0.530/0.528 | -3.857 | 0.542 | -0.700/0.211/0.367/0.398 | 0.788 |
-e2e prompt | 0.399/0.244 | 0.646/0.607/0.624 | -3.363 | 0.575 | -0.194/0.455/0.580/0.605 | 0.982 |
-pipeline prompt | 0.382/0.232 | 0.637/0.597/0.615 | -3.422 | 0.571 | -0.282/0.415/0.554/0.581 | 0.907 |
(a) Direction hDe-En | ||||||
Model | ROUGE-1/L | BERTScore-P/R/F1 | BARTScore | MoverScore | MENLI-/// | DiscoScore |
Supervised Extractive Pipeline | ||||||
MemSum | ||||||
- | 0.315/0.131 | 0.830/0.839/0.834 | -5.320 | 0.814 | -0.648/0.236/0.387/0.418 | 0.301 |
- | 0.311/0.130 | 0.829/0.838/0.833 | -5.350 | 0.814 | -0.652/0.234/0.383/0.413 | 0.291 |
- | 0.269/0.143 | 0.874/0.843/0.858 | -4.914 | 0.824 | -0.638/0.274/0.506/0.552 | 0.605 |
Supervised Abstractive - Baseline CLCTS | ||||||
mLED | ||||||
-Base | 0.321/0.142 | 0.853/0.843/0.848 | -5.275 | 0.823 | -0.605/0.273/0.462/0.499 | 1.653 |
-Historical MT | 0.328/0.143 | 0.853/0.845/0.849 | -5.301 | 0.823 | -0.600/0.276/0.465/0.503 | 1.411 |
0.328/0.144 | 0.854/0.844/0.849 | -5.276 | 0.824 | -0.601/0.276/0.467/0.505 | 1.157 | |
Supervised Abstractive - intermediate finetuning | ||||||
mLED | ||||||
0.316/0.141 | 0.855/0.841/0.848 | -5.253 | 0.826 | -0.613/0.269/0.460/0.498 | 1.500 | |
0.321/0.145 | 0.852/0.841/0.847 | -5.311 | 0.825 | -0.618/0.266/0.453/0.491 | 1.804 | |
-MLS-CLS | 0.328/0.147 | 0.855/0.844/0.850 | -5.250 | 0.826 | -0.586/0.283/0.472/0.510 | 1.050 |
-MLS-CLS-CTS | 0.341/0.153 | 0.855/0.845/0.850 | -5.255 | 0.827 | -0.624/0.268/0.468/0.508 | 0.764 |
Zero-shot Abstractive | ||||||
ChatGPT as summarizer | ||||||
-title prompt | 0.288/0.135 | 0.857/0.840/0.849 | -5.009 | 0.822 | -0.745/0.218/0.444/0.489 | 0.590 |
-e2e prompt | 0.282/0.150 | 0.876/0.843/0.859 | -4.949 | 0.823 | -0.655/0.269/0.509/0.557 | 0.572 |
-pipeline prompt | 0.285/0.147 | 0.874/0.843/0.858 | -4.933 | 0.824 | -0.636/0.274/0.505/0.551 | 0.797 |
(b) Direction hEn-De |
The experiment results are shown in Table 7 for hDe-En (upper) and hEn-De (lower).
Unsupervised extractive pipelines
Finding 1. Slightly Diminishing results from historical text normalization: We observe a decrease in all metrics scores on MemSum pipeline normalized with Norma for both hDe-En and hEn-De. This is contrary to our hypothesis that spelling normalization of historical texts can boost the translation quality for translators originally trained for modern text. Even though Norma performs well at token-level normalization, it fails to consider the context of the words, where in our use case some historical spellings may correspond to multiple modern meanings depending on the context. An example is the word “in” which can function as an adverb in the phrase e.g., “in den Wald” (in the forest) or the accusative case of the word “ihn” (him) in historical texts. A false normalization of such cases results in the meaning change of the entire sentence.
Finding 2. Positive gain from the retrieve-then-summarize pipeline with ChatGPT:
Compared to the extract-then-translate
pipeline, we observe an increase in metric scores for both hDe-En (all metrics except for BARTScore) and hEn-De (except for ROUGE-1). For hDe-En, it is scored as one of the top two models by ROUGE-L, BERTScore
and MENLI variants (in bold). This pipeline is competitive as it not only performs better but expands ChatGPT’s summarization potential for long documents.
Supervised abstractive e2e Finding 1. Mixed results on intermediate task finetuning: For hDe-En, intermediate task finetuning with external sources does not improve scores, and sometimes finetuning with more tasks even deteriorates the outcome according to the metrics. This may be because our CLCTS hDe-En dataset, which dates back to the early 1800s, has substantial linguistic differences to the external sources as discussed in Section 3. For hEn-De, the results show the opposite. We observe higher scores for models trained with more intermediate summarization tasks from external sources. Though, according to MENLI variants and BARTScore, it is not always better to add CTS tasks. This may be because the CTS tasks from HISTSUMM contains documents which have up to 200 years of difference in publication time and again lead to substantial linguistic differences.
Finding 2. Historical translation task improves model performance only mildly:
For hDe-En, the best finetuned model is mLED-Historical MT, where we finetune intermediately with our historical translation task before our target CLCTS task, though according to the metrics, the improvement is very small. Similarly, for the other direction hEn-De, the historical translation task improve model scores but only mildly compared to intermediate tasks. Since the historical translation dataset contains only 201 pairs, a larger size of the historical translation dataset may further improve the results.
Zero-shot abstractive ChatGPT
Finding 1. A close match between pipeline and e2e prompts: For both directions, we observe slightly better scores for summaries generated with e2e prompts which tell ChatGPT to summarize in the target language directly (see Table 6). However, in the following Section 6,
we show that the e2e prompts are prone to invalid outputs.
Finding 2. ChatGPT can output summaries from its memory: For the title prompt experiment, we input zero text from the document but only the title, author and publication year. Slightly surprisingly, for both directions, the resulting scores are mostly comparable to that of the supervised extractive models if not better. ChatGPT is capable of outputting summaries based solely on meta-information for all of our documents. This raises our question on the actual summarization ability of ChatGPT and therefore in the following Section 6,
we design multiple tests to investigate in more detail about ChatGPT involving CLCTS task.
Comparison among all approaches
All the metrics give lower scores to extract-then-translate models (with/without normalization) and ChatGPT title prompt than other models, except for BARTScore, which rates the MemSum pipeline as the highest for hDE-En. However, mixed results are observed from the metrics for determining the best models. BERTScore variants in both directions give higher scores to ChatGPT-related models (excluding title prompt). MENLI variants agree with this fully for hDe-En, and partially agree for hEn-De. The same is also observed from MoverScore and ROUGE-L. BARTScore, however, contradicts with them and scores ChatGPT models higher for hEn-De, though only partly for hDe-En. Moreover, for hEn-De, ChatGPT is no longer dominating. However, the study from Goyal, Li, and Durrett (2022) also shows that current metrics fail to evaluate zero-shot model output properly. To better understand the performance of models and the metrics, we also consider human evaluation.
5.2 Human and ChatGPT evaluation
Table 8 shows the average Spearman’s ranking correlation coefficient among the annotators for coherence, consistency, fluency, and relevance. We obtain moderate agreement for relevance in both directions and moderate agreement for coherence and consistency in hDe-En. The annotation agreement for hEn-De is slightly worse at a weak-moderate level except for relevance. One exception is fluency in hDe-En, where the correlation is poor.
Models | Coh. | Con. | Flu. | Rel. |
hDe-En | 0.537 | 0.612 | 0.176 | 0.600 |
hEn-De | 0.353 | 0.493 | 0.433 | 0.605 |
Model | Coherence | Consistency | Fluency | Relevance |
Supervised Extractive | ||||
MemSum | ||||
- | 2.25/1.77 | 1.75/1.74 | 1.75/1.62 | 2.25/1.46 |
- | 2.25/1.71 | 1.50/1.39 | 2.75/1.63 | 1.00/1.18 |
- | 3.75/3.81 | 2.75/4.14 | 4.00/3.84 | 3.00/3.97 |
Supervised Abstractive - Baseline CLCTS | ||||
mLED | ||||
-Base | 2.10/1.88 | 2.08/1.88 | 2.78/2.00 | 2.22/1.62 |
-Historical MT | 2.76/2.00 | 2.44/1.80 | 2.71/2.00 | 2.54/1.80 |
- | 2.38/1.50 | 2.21/1.17 | 2.49/1.83 | 2.30/1.17 |
Supervised Abstractive - intermediate finetuning | ||||
mLED | ||||
2.42/2.00 | 2.00/2.00 | 2.92/2.00 | 2.67/1.33 | |
1.94/1.25 | 1.88/1.75 | 2.69/1.25 | 2.25/1.25 | |
-MLS-CLS | 1.88/1.50 | 1.50/1.25 | 2.81/1.25 | 2.06/1.00 |
-MLS-CLS-CTS | 2.09/1.71 | 1.89/1.29 | 2.43/1.71 | 2.04/1.14 |
Zero-shot Abstractive - ChatGPT | ||||
ChatGPT as summarizer | ||||
-title prompt | 2.88/2.75 | 3.25/2.25 | 4.75/3.25 | 2.31/2.25 |
-e2e prompt | 3.80/3.79 | 3.50/4.23 | 4.00/3.87 | 4.00/4.08 |
-pipeline prompt | 4.50/4.25 | 4.38/4.50 | 5.00/4.00 | 4.25/4.25 |
(a) Direction hDe-En | ||||
Model | Coherence | Consistency | Fluency | Relevance |
Supervised Extractive | ||||
MemSum | ||||
- | 2.30/1.75 | 1.60/1.43 | 1.80/1.75 | 2.00/1.10 |
- | 2.30/1.68 | 1.40/1.40 | 2.60/1.57 | 1.00/1.00 |
- | 3.88/3.44 | 3.00/3.56 | 3.25/3.56 | 3.00/3.30 |
Supervised Abstractive - Baseline CLCTS | ||||
mLED | ||||
-Base | 2.01/1.29 | 1.50/1.00 | 2.30/1.57 | 1.60/1.00 |
-Historical MT | 1.89/1.67 | 1.79/1.17 | 2.82/1.67 | 1.79/1.00 |
- | 2.00/1.00 | 1.62/1.00 | 1.62/1.00 | 1.62/1.00 |
Supervised Abstractive - intermediate finetuning | ||||
mLED | ||||
1.99/1.50 | 1.79/1.33 | 2.40/1.50 | 1.96/1.17 | |
1.88/1.29 | 1.80/1.07 | 2.38/1.29 | 1.75/1.00 | |
-MLS-CLS | 1.76/1.33 | 1.67/1.33 | 2.38/1.50 | 1.97/1.00 |
-MLS-CLS-CTS | 2.38/1.50 | 2.00/1.00 | 1.88/1.50 | 1.75/1.00 |
Zero-shot Abstractive - ChatGPT | ||||
ChatGPT as summarizer | ||||
-title prompt | 3.12/2.50 | 2.12/2.25 | 2.81/3.00 | 2.12/2.00 |
-e2e prompt | 3.75/3.21 | 3.13/3.18 | 3.25/3.36 | 3.13/2.85 |
-pipeline prompt | 3.06/3.25 | 3.38/3.25 | 3.06/3.50 | 3.38/3.25 |
(b) Direction hEn-De | ||||
Models | Coh. | Con. | Flu. | Rel. |
hDe-En | 0.490 | 0.569 | 0.597 | 0.577 |
hEn-De | 0.320 | 0.377 | 0.392 | 0.455 |
(c) Annotation agreement for human and ChatGPT |
We report the mean value of human and ChatGPT annotations for each model in Table 9. The best scores for supervised abstractive models in both directions range between 2 to 3 by human annotators and only above 1 by ChatGPT, which indicates the quality of outputs is bad to mediocre. However, the highest scores for ChatGPT as a summarizer are over 4 (very good) for hDe-En and 3 to 4 (mediocre to good) for hEn-De. Overall, the quality of outputs for hEn-De is lower than that for hDe-En. This is also expected since the length of historical English documents is much longer with 9643.2 words on average which results in more information loss due to text truncation because of computational limitations and the cross-lingual embedding distance from Table 3 also suggests more difficulty of hEn-De compared to hDe-En.
For both directions, humans give higher scores to summaries generated by ChatGPT, especially ChatGPT pipeline prompt for hDe-En with an increase of a great margin. Unlike evaluation metrics, humans prefer pipeline prompt output better than e2e prompt output for hDe-En and for hEn-De in coherence and relevance. A similar finding also holds for comparing finetuned models, where humans give higher scores for the model finetuned with historical translation similar to evaluation metrics for hDe-En. The ratings for hEn-De differs slightly
where human ratings agree with metric evaluation results in terms of coherence and consistency. Since the overall quality of generated summaries is bad for hEn-De, the comparisons among outputs may be more difficult for both humans and evaluation metrics.
Mixed results for ChatGPT as an evaluator.
Table 9 (c) shows the annotation agreement between humans and ChatGPT annotation. For hDe-En, the agreement is still moderate but weak for hEn-De overall. Similarly to the findings from Chiang and yi Lee (2023), we observe lower Likert scores from ChatGPT evaluation compared to human evaluation except for summaries generated by ChatGPT itself. However, the output for hEn-De is more concerning where we observe many ratings of score 1, especially for relevance. Another point worth noticing is that even though in our instruction, we allow for 0.5 increments, we do not observe such ratings from ChatGPT output.
To sum up, ChatGPT as an evaluator (i) is prone to lower scores than humans 101010We include one sample annotation from the gold standard annotations during instruction.; (ii) favours its own output since we observe similar (even higher) scores from ChatGPT evaluation compared to human evaluation for ChatGPT summarization outputs; (iii) cannot distinguish summaries of low qualities (hEn-De); (iv) does not fully follow the instruction of 0.5 score increments.
6 Discussion and analysis
6.1 Analysis on ChatGPT summarization
6.1.1 The effect of different prompts and target language
A case study with ChatGPT-e2e vs. ChatGPT-pipeline. Though in Section 5 we observe better metric and human annotation results from ChatGPT e2e prompt, we notice that ChatGPT under this prompt is prone to output summaries in the wrong target language (invalid output). In such cases, we query the summary with the same prompt until a valid output is returned. We calculate the occurrences of invalid outputs after 2 rounds of queries and report them in Table 10, where we observe that using e2e prompts has a higher chance to yield invalid outputs. What is also worth noticing is when the output is expected in German, we observe more invalid cases among all three prompts which means ChatGPT tend to output English texts when our expected language is in German.
Model | Trg.lang | Invalid/Obs. |
ChatGPT-title | De | 1/328 |
ChatGPT-pipeline | De | 34/328 |
ChatGPT-e2e | De | 57/328 |
ChatGPT-title | En | 0/289 |
ChatGPT-pipeline | En | 2/289 |
ChatGPT-e2e | En | 8/289 |
6.1.2 ChatGPT for historical text translation
[…]
6.1.3 ChatGPT knows it all?
Since our corpus-building process is highly dependent on Wikipedia (and popular stories presumably discussed on other places on the web), we want to know to what extent ChatGPT has an unfair advantage. We ask ChatGPT to output summaries based on the author, title and publication year. ChatGPT reports summaries to all requests in both language directions, though the output may contain irrelevant information besides the actual summarization.111111E.g. “The story emphasizes the power of kindness and the importance of inner beauty. It also highlights the value of perseverance and the rewards that can come from staying true to oneself.”
However, despite the overall outstanding performance of ChatGPT summarization according to both humans, ChatGPT evaluation and automatic evaluation metrics, ChatGPT as a summarizer can also be prone to hallucination like other NLG systems Bang et al. (2023); Ji et al. (2023). For example, in the summary of the famous story Cinderella, we notice the golden slippers in the original document become glass slippers which is the best-known version worldwide but is actually incorrect based on the input text.
Adversarial attacks for ChatGPT
Knowing that ChatGPT has specific prior knowledge of our CLCTS task, we take a step further and design three experiments inspired by the work from Chen and Eger (2023) where we gradually reduce the amount of specific prior knowledge in the input documents (from seen documents to pseudo-unseen documents to unseen documents). We aim to answer (i) does ChatGPT summarize truly based on the input text given its specific prior knowledge? (ii) Can
ChatGPT summarize against its specific prior knowledge? (iii) How well can ChatGPT summarize facing unseen documents?
For question (i), we design sentence omission task. By dropping an increasing amount of sentences from the original documents, we control the information reduction of ChatGPT input for summarization. We randomly select 11 historical English documents from our CLCTS corpus containing 100-150 sentences and construct new documents by dropping sentences of different percentages. We then query summaries of the constructed documents through ChatGPT. To quantify the information change, we use the summary of the original documents (generated by ChatGPT) as the reference, and evaluate generated summaries from documents after omission using automatic evaluation metrics.121212We set the temperature to 0 since our goal is not to obtain the best summaries but to test ChatGPT summarization performance under gradual information loss. We average the resulting metric scores after scaling and use it as our similarity measure. The result is shown in Figure 4. We observe an overall downward trend of similarity score of both CTS and CLCTS which reflects a certain level of information loss accordingly. This indicates that ChatGPT summarizes according to the given information in general which contradicts our hypothesis that ChatGPT will be affected by its prior knowledge where the curve would be more flat.
For question (ii), we first construct pseudo-unseen documents by entity swap, negation, and omission.131313For each attack, we construct 10 documents. For each document, we query the summaries 3 times at temperatures 0, 0.7 and 1. In total, we exclude summaries in the wrong target language and in the end, we have 257 valid generated summaries. Entity swap: We swap the entities from the original story with new entities. For example, in the story of Cinderella, we switch the hazel tree into cactus, or (to make it funny) we replace the prince with a watermelon (to ensure coherence, we also add background such as “They live in the Fruit Kingdom.”). Negation: We alter the plot completely to make a negated story. For example, in the story of Cinderella, the stepmother and stepsisters are mistreated by the rich man’s evil daughter and the prince eventually marries one of the stepsisters. Omission: For omission, we omit important plots. For example, Cinderella neither attends the ball nor marries the prince. We make sure each pseudo-unseen document only contains one kind of attack. We manually annotate the resulting output where we focus on if the attack is completed accurately or not (e.g., if during negation, Cinderella is mistreated again by the stepmother, then the task failed). We compute the accuracy and report it in Table 11.
Attack | CTS | CLCTS |
Omission | 0.74 | 0.63 |
Entity Swap | 0.80 | 0.71 |
Negation | 0.69 | 0.53 |
We observe from the results that (a) ChatGPT can better handle omission and entity swap than negation. This may be because compared to omission and entity swap, negation attacks alter the underlying logical connections of the story and thus requires more reasoning to complete this attack under the assumption that ChatGPT is affected by its explicit knowledge during summarization. (b) Higher accuracy for CLCTS than CTS task. For all three attacks, we observe a higher accuracy of completing the task for CTS. This is understandable since CLCTS also involves summarization cross-lingually which is more difficult.
To further investigate question (iii), we create 5 unseen documents from fiction chapters whose last updates are in year 2022 or later141414We include chatpters from 3 Chinese novels and 2 English novels.. For modern fiction chapters, we translate and rewrite the stories in historical languages and write reference summaries in modern languages. Both documents and summaries are proofread by a co-author of this work. Similar to Section 5, we evaluate the resulting summaries by three human annotators from four perspectives. The results are shown in Table 12. We notice that the overall rating for CTS and CLCTS ranges between 3 to 4 indicating that ChatGPT summarizes unseen documents at a mediocre level. This rating is similar to the results from our previous experiment using CLCTS corpus (we list the result of the same direction in Table 5).
ChatGPT can outperform finetuned models in terms of CLCTS tasks in our experiment. However, it does not magically solve all the issues as per our test, ChatGPT is better at omission and entity swap than negating against its prior knowledge and it generates summaries of mediocre quality from unseen source documents.
Models | Coh. | Con. | Flu. | Rel. |
unseen CTS hEn-En | 3.38 | 3.00 | 3.63 | 3.38 |
unseen CLCTS hEn-De | 3.40 | 3.17 | 3.78 | 3.17 |
our CLCTS hEn-De | 3.06 | 3.38 | 3.06 | 3.38 |
(a) Direction hDe-En
(b) Direction hEn-De
6.2 How well do different metrics perform at segment level?
Figure 5 shows Spearman’s ranking correlation between human annotation and evaluation metrics at segment level.
Decent correlation from BERTScore variants but worse for German For hDe-En, we observe a weak to moderate level of correlation with human annotation for BERTScore, and . However, the correlations for MENLI variants are negative. BARTScore correlates weakly in coherence and relevance. ROUGE-1 and ROUGE-L have low correlations with human annotation and the same is also true for MoverScore and DiscoScore. The situation is worse for hEn-De where we observe a very low level of correlation in general and less significant correlations from BERTScore variants according to the Student’s t-test but ROUGE-L correlates moderately in hEn-De in consistency and relevance. Overall, BERTScore variants correlate slightly better than other metrics with humans in hDe-En for coherence (0.49), consistency (0.51) and relevance (0.5), and in hEn-De for consistency (0.41) and relevance (0.26).
[…]
Additionally, we compute the correlation between evaluation metrics and human annotation among ChatGPT outputs only. Surprisingly, the resulting correlation is again moderate and even strong for BERTScore-P and DiscoScore (in consistency & relevance for hDe-En) though not significant according to a two-tailed Student’s t-test. However, this fact is still interesting since it may suggest that evaluation metrics work better when comparing outputs from similar training paradigms but worse when comparing between them (e.g., finetuned and prompt-based models).
7 Concluding remarks
In this work, we build the first CLCTS corpus and a historical translation dataset with additional information for future explorations. We study the characteristics of this corpus both alone and in comparison to other summarization corpora and examine the effectiveness of pipeline models, popular e2e transformer-based abstractive models leveraging intermediate task finetuning and ChatGPT for the CLCTS task. Even though these methods have been shown effective for single-task settings such as CLS or long document summarization, they fail to deliver good summaries in our combined task setting (cross-lingual, cross-temporal and long document). We show the potential of ChatGPT in CLCTS both as a summarizer and an evaluator coupled with different attacks to unveil its strength and weakness. Additionally, we find a moderate correlation of BERTScore variants with human annotation (slightly worse for German) and an overall improvement of correlation after excluding ChatGPT outputs at segment level.
Future work can address the limitations revealed by our experiments, including experiments with other models of great potential (e.g., unlimiformer Bertsch et al. (2023)), with other training frameworks (e.g., pertaining under multi-task learning) or with more historical data focusing on other languages (also more distant ones) and tasks (e.g., non-parallel historical corpora for pretraining and PAR3 cross-temporal dataset for machine translation Thai et al. (2022)). Furthermore, future work can focus on extending the size and diversity of our CLCTS datasets.
8 Acknowledgments
We thank all annotators for their hard work and time. Ran Zhang is financed by the BMBF project “Metrics4NLG”. Steffen Eger is financed by DFG Heisenberg grant EG 375/5–1.