### v3

# Cross-lingual Cross-temporal Summarization: Dataset, Models, Evaluation

Abstract
While summarization has been extensively researched in natural language processing (NLP), cross-lingual cross-temporal summarization (CLCTS) is a largely unexplored area that has the potential to improve cross-cultural accessibility and understanding. This paper comprehensively addresses the CLCTS task, including dataset creation, modeling, and evaluation. We (1) build the first CLCTS corpus with 328 instances for hDe-En (extended version with 455 instances) and 289 for hEn-De (extended version with 501 instances), leveraging historical fiction texts and Wikipedia summaries in English and German; (2) examine the effectiveness of popular transformer end-to-end models with different intermediate finetuning tasks; (3) explore the potential of GPT-3.5 as a summarizer; (4) report evaluations from humans, GPT-4, and several recent automatic evaluation metrics. Our results indicate that intermediate task finetuned end-to-end models generate bad to moderate quality summaries while GPT-3.5, as a zero-shot summarizer, provides moderate to good quality outputs. GPT-3.5 also seems very adept at normalizing historical text. To assess data contamination in GPT-3.5, we design an adversarial attack scheme in which we find that GPT-3.5 performs slightly worse for unseen source documents compared to seen documents. Moreover, it sometimes hallucinates when the source sentences are inverted against its prior knowledge with a summarization accuracy of 0.67 for plot omission, 0.71 for entity swap, and 0.53 for plot negation. Overall, our regression results of model performances suggest that longer, older, and more complex source texts (all of which are more characteristic for historical language variants) are harder to summarize for all models, indicating the difficulty of the CLCTS task. Regarding evaluation, we observe that both GPT-4 and BERTScore correlate moderately with human evaluations, implicating great potential for future improvement.
1 Introduction
Summarization is a key task in natural language processing (NLP), especially in an age of information overload. The classical approach for summarization is to summarize documents (e.g., news reports) in the same language Zhang et al. (2020a); Liu et al. (2022); Ravaut, Joty, and Chen (2022) or, less prominently, in different languages than the source (cross-lingual summarization; CLS) Ladhak et al. (2020); Cao, Liu, and Wan (2020); Bai, Gao, and Huang (2021); Liang et al. (2022). Tasks such as the summarization of historical texts to modern languages (cross-temporal summarization; CTS) are scarcely explored, in contrast. Such summarizers are beneficial not only for historical researchers and students but also for laypeople with an interest in historical information sources.
In this work, we go one step further and explore cross-lingual cross-temporal summarization (CLCTS), in which the goal is to summarize a historical document in a different modern language. See Table 1 for an example. Such cross-lingual and cross-temporal summarizers would potentially tremendously facilitate cross-cultural accessibility, information sharing, and understanding. CLCTS is a valuable area of research that expands upon the recently popular task of literary translation with summarization. Both tasks share cultural importance and present unique complexities due to the intricate nature of the creative work involved Karpinska and Iyyer (2023); Thai et al. (2022). CLCTS is not limited to literary works alone: it can also assist in summarizing non-literature such as historical newspapers, reports, textbooks, etc. For example, CLCTS models can assist in creating (missing) Wikipedia pages for historical works. While the translation of literature may require language that is faithful to the original source for a ‘close reader’, CLCTS could potentially be better suited for a ‘distant reader’, and thus may require less sophisticated output text. Moreover, modern summaries can facilitate the classification of literary works, for example, genre classification Saraswat and Srishti (2022); Agarwal, Vijay et al. (2021). Firstly, summaries are more condensed than (historical) source texts and, therefore, are much easier to model. Secondly, modern texts are better suited for models that operate with modern data such as SBERT Reimers and Gurevych (2020).
Title | An Inhabitant of Carcosa |
Author | Ambrose Bierce |
Year | 1886 |
Source | The Collected Works of Ambrose Bierce, Volume 3 |
Text | “For there be divers sorts of death—some wherein the body remaineth; and in some it vanisheth quite away with the spirit … In one kind of death the spirit also dieth, and this it hath been known to do while yet the body was in vigor for many years. […]” |
Summary (German) | “Ein Mann aus der Stadt Carcosa, der über die Worte des Philosophen Hali über die Natur des Todes nachdenkt, wandert durch eine ihm unbekannte Wildnis. […]” |
CLCTS is difficult and challenging for the following reasons: (i) languages change over time along multiple dimensions, including syntax Juzek, Krielke, and Teich (2020); Lei and Wen (2020), semantics Hamilton, Leskovec, and Jurafsky (2016); Giulianelli, Del Tredici, and Fernández (2020), lexical choice and morphology Gibson et al. (2019); Joseph (2017). (ii) Historical documents are often longer, not only because of tendencies of simplification/text length reduction over time Lewis (1894); Sherman (1893); Rudnicka (2018), but also because of the genre of historical text, which predominantly includes literary work Wang and Liu (2017); Zhu, Liu, and Pang (2022). (iii) The process can be thought of as consisting of several subtasks, previously considered independently in NLP research: machine translation (MT), summarization, and historical (spelling) normalization Bollmann and Søgaard (2016); Eger, vor der Brück, and Mehler (2016).
In this work, we consider the CLCTS problem comprehensively, including dataset creation, modeling, and evaluation.111Our code and data are available at https://github.com/zhangr2021/CLCTS. We start with building the first CLCTS dataset. Our corpus contains stories and plays (as prime artefacts of historical cultures), which requires the models to handle long documents; we assemble modern summaries by querying Wikipedia.
In terms of modeling, we consider three approaches: (1) extractive summarizers Gu, Ash, and Hahnloser (2022) whose advantage is that they can in principle deal with unlimited input document lengths but which are limited in that they can only copy out sentences from the original source (which is a problem, e.g., when the source is first-person narrative but the summary should be third-person perspective). (2) Abstractive summarizers can avoid the issue above but typically have severe limitations in input lengths. Even models exclusively designed for long document summarization by utilizing efficient variants of attention mechanism such as Longformer Encoder-Decoder (LED) Beltagy, Peters, and Cohan (2020) and Bigbird Zaheer et al. (2020) limit the max input tokens to 16,384 and 4,096, respectively. (3) GPT-3 Brown et al. (2020) and its extensions such as ChatGPT are also strong candidates for our task Goyal, Li, and Durrett (2022a) as they have so spectacularly impacted research in a multitude of scenarios Leiter et al. (2024), including summarization Yang et al. (2023); Wang et al. (2023); Zhang, Liu, and Zhang (2023). Consequently, we also include ChatGPT (GPT-3.5 and GPT-4) in our analysis.
In terms of evaluation, we consider three evaluation methods: (1) automatic evaluation, where we report multiple evaluation metrics additional to the commonly used variants of ROUGE Lin (2004) in summarization tasks Liang et al. (2022); Bai, Gao, and Huang (2021); Cao, Liu, and Wan (2020); Peng et al. (2021) since the performance of individual metrics may vary across datasets, challenging their reliability Bhandari et al. (2020); Fabbri et al. (2021); (2) human evaluation, which reflects the actual quality of summaries according to human judgments and functions as a source of reliability measurement for automatic evaluation metrics; (3) ChatGPT evaluation, where we examine the potential of ChatGPT as an alternative to human annotators given the same instruction considering the high cost of human evaluation Gao et al. (2023) and issues of reproducibility Belz et al. (2021); Belz, Thomson, and Reiter (2023); Chen, Belouadi, and Eger (2022). Our contributions are:
To our best knowledge, we build the first CLCTS corpus, leveraging historical fiction texts and Wikipedia summaries.
We examine the effectiveness of popular transformer end-to-end (e2e) models with different tasks for intermediate task finetuning Chang and Lu (2021).
We explore how ChatGPT performs for CLCTS.
We provide human, GPT-4, and automatic evaluation for the task, showing that our intermediate task finetuned e2e models generate bad to moderate quality summaries; GPT-3.5 as a summarizer provides moderate to good quality outputs (without any finetuning); GPT-4 as an evaluator correlates with human evaluations to a moderate level.
2 Related work
Historical text normalization Historical text normalization, the process of mapping non-standard word tokens to their modern standard forms, is an important technique for analyzing historical texts. Model designs include (1) corpus-based lexical substitution as an effective component for normalization systems Rayson, Archer, and Smith (2005); Baron and Rayson (2008); Bollmann (2012); (2) rule-based models, such as phonological rules Porta, Sancho, and Gómez (2013) and edit distance measures Adesam, Ahlberg, and Bouma (2012); Bollmann (2012); (3) statistical models which aim to maximize the probability of contemporary word form given its historical form Pettersson (2016); (4) neural network (NN) models which utilize NN architectures such as encoder-decoder LSTMs Bollmann, Bingel, and Søgaard (2017); Robertson and Goldwater (2018) or RNNs Makarov and Clematide (2020). Worth noting is that nearly all existing historical normalizers operate on words in isolation, except for the semi-supervised contextual normalizer from Makarov and Clematide (2020). In our study, we find that ChatGPT (GPT-3.5) has the potential to serve as a (much better) context-aware historical normalizer.
[…]
Since we focus on various summarizations instead of translations and long documents instead of paragraphs, the translation dataset is not within the scope of our current study.
Cross-lingual summarization
[…]
Long document summarization is another area of our concern.
[…]
We build our model with mBART Tang et al. (2020) pretrained for multilingual MT tasks and combine with LED attention to enable efficient processing for long sequences. Additionally, we build another retrieve-then-summarize pipeline where we retrieve sentences with an extractive summarizer Gu, Ash, and Hahnloser (2022) and then summarize cross-lingually using ChatGPT.
ChatGPT for summarization Despite ChatGPT’s young age, there is already a wealth of research that explores it for various NLP tasks, including summarization Yang et al. (2023) and summarization evaluation Chiang and Lee (2023); Gao et al. (2023). Disagreement exists concerning ChatGPT’s performance.
Most studies focus on MLS tasks.
[…]
They also claim that ChatGPT can perform better with pipeline-like prompt lines
in an interactive manner. It is also worth noticing that all works mentioned above utilize small samples of 50 to 100 random samples (out of 3,000 to 60k observations) from each dataset in their experiments.
[…]
3 Dataset
In this section, we introduce our CLCTS corpus and additional sources we utilize for our experiments. We build our own corpus which covers English and German in both directions. We utilize multiple monolingual and cross-lingual summarization datasets as additional resources during our intermediate task finetuning process.
3.1 CLCTS Corpus
Manually creating summaries is very time-consuming and requires expert knowledge Peng et al. (2021). Thus, we use fairy tales, short stories, and plays where summaries are available in Wikipedia articles. In each language direction, English and German, the dataset contains summary pairs from historical documents in one language to modern summaries in the other language.
Dataset collection We collect historical documents mainly from three different sources.555Wikisource and Project Gutenberg contain both German and English sources, while Deutsches Textarchiv contains German sources only. The links to the sources are collected in Section 9.1 (appendix).
Deutsches Textarchiv (DTA; German text archive): This is a basic stock of German-language texts focusing on the early 16th to early 20th centuries.
Wikisource: This is a multilingual online digital library of free-textual content containing historical books in text format and original format stored as images. Some characters are changed from the original historical format to modern digital formats. Proofreading is performed for the texts by users on Wikisource following the proofreading guide.
Project Gutenberg: This is a cultural digital archive with over 60,000 digitized books, including historical texts. It focuses on older works with expired U.S. copyrights.
Our workflow for building the corpus consists of two stages: (1) meta-information collection & historical text extraction and (2) summary collection & translation. In stage (1), we collect meta-information of historical documents using the “Beautifulsoup” library. We extract information such as titles, web links, authors, years, and documents together with information from Wikipedia lists of German fairy tales and English short stories. For documents presented in several sources, the oldest version or the version with historical spelling is selected. For stage (2), we utilize the collected information to match the historical documents with their corresponding summaries from Wikipedia. The historical spelling of titles is converted to modern spelling to match the titles in Wikipedia. Figure 1 illustrates the process of data collection. If no original human correspondent summaries are found in the other language, we translate the summaries into the target language using DeepL (we translate 144 out of 328 German summaries to English and 108 out of 289 English summaries to German).666DeepL API: https://www.deepl.com/pro-api An example of our collected dataset is shown in Table 1.
In addition to the summarization datasets, we create a historical German-English translation dataset with 201 translation pairs. This dataset is later utilized in intermediate task finetuning. We retrieve the German documents from the DTA corpus and the historical English from Wikisource. The documents are translated by Margaret Raine Hunt (1831–1912). An example of the translation dataset is given in Table 17 (appendix).
We use the abbreviations hDe and hEn for historical German and historical English texts to distinguish them from modern texts De or En. Thus, we refer to the historical German to modern English summarization dataset as “hDe-En” and the historical English to modern German summarization dataset as “hEn-De”. Similarly, monolingual CTS datasets are called “hDe-De” or “hEn-En”. Names of other modern datasets follow the same rule.
3.2 External datasets
To enrich our finetuning data resources, we utilize external summarization datasets: (1) HISTSUMM Peng et al. (2021) where we also translate the summaries into English using DeepL to enable cross-lingual finetuning. (2) We further use Wikilingua for MLS and CLS tasks Ladhak et al. (2020). Wikilingua is a multilingual summarization dataset that contains 18 languages from WikiHow (an online Wiki-style publication featuring how-to articles), including English and German. The English dataset contains 141,457 document-summary pairs. For finetuning our models, we randomly select 60% of the pairs (85,874) and we use all of the German dataset (58,341). Wikilingua German entries contain the URL and the section name of the English articles. Thus, cross-lingual summary pairs can be retrieved using this information and we match 20,103 cross-lingual summary pairs in both directions. (3) We also utilize the CNN/Daily Mail dataset. This is a monolingual English summarization dataset containing 311,971 document-summary pairs, including news articles written by CNN and Daily Mail journalists, and summaries from the highlight of the article written by the article author. We experiment with a subset of 86,133 text-summary pairs. (4) We include MLSUM (Scialom et al., 2020), which is a multilingual summarization dataset containing over 1.5 million document-summary pairs collected from online newspapers such as Süddeutsche Zeitung (German). We use a subset of 66,226 pairs from the German dataset.
3.3 Dataset Statistics
Dataset | Size | Mean Length | Compression | Genre | Task | |||
Doc. | Summ. | |||||||
CLCTS Corpus | ||||||||
CLCTS hEn-De | 289 | 9,643.3 | 414.0 | 16.3 | 19.2 | 23.3 | fiction | CLS, CTS |
CLCTS hDe-En | 328 | 1,398.4 | 324.3 | 25.4 | 18.6 | 4.3 | fiction | CLS, CTS |
CLCTS hEn-En | 276 | 9,847.2 | 488.1 | 16.0 | 20.2 | 20.2 | fiction | MLS, CTS |
CLCTS hDe-De | 328 | 1,398.4 | 217.4 | 25.4 | 16.6 | 6.4 | fiction | MLS, CTS |
External Corpus | ||||||||
HISTSUMM hDe-De | 100 | 268.1 | 18.1 | 36.3 | 12.4 | 14.8 | news | MLS, CTS |
Wikilingua En-En | 85,874 | 425.8 | 39.6 | 16.5 | 6.5 | 10.8 | how-to-guide | MLS |
Wikilingua De-De | 58,341 | 429.8 | 42.2 | 15.8 | 6.9 | 10.2 | how-to-guide | MLS |
Wikilingua De-En | 20,103 | 438.2 | 38.8 | 15.8 | 6.5 | 11.3 | how-to-guide | CLS |
Wikilingua En-De | 20,103 | 451.7 | 42.7 | 16.5 | 6.9 | 10.6 | how-to-guide | CLS |
CNN/Daily Mail En-En | 86,133 | 786.7 | 55.1 | 20.5 | 14.5 | 14.3 | news | MLS |
MLSUM De-De | 66,226 | 570.3 | 30.4 | 17.8 | 12.7 | 18.8 | news | MLS |
In Table 2, we report statistics of our corpus and other existing summarization datasets. We observe three trends comparing all datasets:
Much fewer instances in Historical datasets. The CLCTS hEn-De dataset consists of 289 historical short stories and plays, and the CLCTS hDe-En dataset consists of 328 historical fairy tales. 777We limit the maximum length of the document to 16k. Since we collect summaries from Wikipedia, the total number of matched document-summary pairs is limited. Stories with less popularity are less likely to be found on Wikipedia than well-known ones. The size of HISTSUMM dataset is even smaller, with only 100 historical news articles. In contrast, modern summarization datasets contain several thousand summarization pairs, which is more than 100 times the size of historical datasets. The largest dataset CNN/Daily Mail contains over 300k summarization pairs.
Longer documents and summaries for CLCTS corpus. The CLCTS hEn-De dataset has more than 9,643 tokens per document and 414 tokens per summary on average and CLCTS hDe-En has 1,398 per document and 324 per summary. For external datasets, the average length is below 800 tokens per document and below 60 tokens per summary. This is because the CLCTS corpus includes historical documents from the genre fiction, which are naturally longer than news and WikiHow (how-to guide) due to genre differences (Rudnicka, 2018) and higher verbosity of more historical languages.
Larger information loss during summarization for news genre and CLCTS hEn-De measured by compression ratio. News documents have a higher compression ratio (the mean length of document divided by the mean length of summary) compared to documents from other genres, except for CLCTS hEn-De, which has the highest compression ratio among all datasets. The CLCTS hEn-De dataset contains short stories and plays that are much longer than fairytales in CLCTS hDe-En and documents from all other genres as discussed above. For the CLCTS corpus, the compression ratio differs substantially because all summaries of the CLCTS corpus are collected from Wikipedia with the average length of summaries ranging from 200 to 400 tokens regardless of the original document length.
To obtain more information on document-summary pairs, we quantify the divergence of sentence semantics between document and summary by computing the mean sentence-level cosine similarity for all datasets (Vyas, Niu, and Carpuat, 2018). We obtain the sentence embedding via multilingual SBERT (Reimers and Gurevych, 2020). We first compute the average similarity of each source to each summary sentence to obtain document level similarity and then average over all documents to obtain dataset level similarity score. The results are shown in Table 3. It reveals that cross-lingual sentence semantics between document and summary is more diverged than that of monolingual datasets and among all, both CLCTS hDe-En and hEn-De (which are our target tasks) have lower embedding similarity compared to other datasets. This again indicates the difficulty of cross-lingual and cross-temporal summarization.
Monolingual | Cross-lingual | ||
CLCTS hDe-De | 0.38 | CLCTS hDe-En | 0.33 |
CLCTS hEn-En | 0.27 | CLCTS hEn-De | 0.25 |
HISTSUMM hDe-De | 0.39 | - | - |
CNN/Daily Mail En-En | 0.32 | - | - |
MLSUM De-De | 0.37 | - | - |
Wikilingua De-De | 0.46 | Wikilingua De-En | 0.42 |
Wikilingua En-En | 0.42 | Wikilingua En-De | 0.42 |
We include an analysis of mean sentence length in the next section where we focus on a detailed analysis of historical documents and their comparison to modern documents.
3.4 Historical Language Divergence
Languages evolve due to social, cultural, and linguistic pressure and this poses challenges to the summarization of historical texts. In this section, we firstly provide an overview of the publication time of the historical documents from HISTSUMM and CLCTS and secondly, analyze the divergence of historical language from syntax (including the length of sentences and dependency distances) and lexicon perspectives.
Publication time of historical documents
The distribution of publication year is shown in Figure 2. HISTSUMM documents (in red) distribute evenly between the period of 1650–1800. CLCTS hEn-De (in gray) contains more documents from the late 1800s and CLCTS hDe-En (in green) from the early and mid 1800s. The documents from CLCTS hEn-De which date back to the 1600s are plays from Shakespeare. For both CLCTS datasets, the majority of historical documents come from the 1800-1900s and comparatively HISTSUMM has overall older documents than our CLCTS corpus.
Divergence in syntax Length of Sentences The analysis in Table 2 (column 5) reveals that historical German documents in HISTSUMM and CLCTS corpus contain longer sentences compared to other datasets, where HISTSUMM hDe-De has a mean sentence length of 36.3 tokens and CLCTS hDe-En/hDe-De dataset has 25.4 tokens per sentence while modern German documents from Wikilingua and MLSUM have maximum 20 tokens per sentence on average. This coincides with the finding of decreasing sentence length over time (Rudnicka, 2018), considering that HISTSUMM documents are the oldest. Slightly surprisingly, this is not true for the English documents, where the mean sentence length of historical English documents (CLCTS hEn-De/hEn-En) is almost the same as Wikilingua and is even shorter than documents from the news genre CNN/ Daily Mail. This hints at a genre effect, besides the temporal effect: fiction, especially plays, may have shorter sentences than news and other non-fiction genres due to the higher occurrence of dialogues and two-word sentences (Rudnicka, 2018). In our CLCTS corpus, unlike CLCTS hDe-En/hDe-De containing only fairytales, historical English documents contain plays. When we exclude plays from Shakespeare, the mean sentence length of the CLCTS hEn-De document increases by 4 tokens per sentence to 20.3, which is closer to the value of CNN/Daily Mail. The factual summaries from CLCTS hEn-De/hEn-En in column 6 also contain longer sentences than their fictional documents for similar reasons.
Dependency Distances To further quantify syntactic changes, we analyze the dependency parsing trees for sentences from documents in all datasets and compute the mean dependency distance (MDD) over the entire corpus Liu (2008); Liu, Xu, and Liang (2017). This measure can reflect the syntactic complexity of sentences. We use stanza (Qi et al., 2020) for parsing the sentences. We plot the MDD in Figure 3 where we observe a downward trend overall. The first four documents from the left are in German and the last three documents are in English (we use dataset plus language, e.g., CLCTS hDe to represent the documents here). The downward trending line reveals an overall higher MDD of German sentences than that of English sentences. If we take a closer look by language, we observe: (1) For German (the first four points from the left), historical sentences (in red star) have a higher value for MDD than that of the modern German text, which suggests more syntactic complexity of historical documents. We especially notice a higher MDD from CLCTS hDe documents (fiction) than that from HISTSUMM hDe (news), which coincides with the finding of longer dependencies in the imaginative genre than those in the informative genre from Wang and Liu (2017). The modern sentences from MLSUM are also more complex syntactically compared to Wikilingua. (2) For English (top-3 right items), this is not entirely true. News documents from CNN/Daily Mail have again a longer MDD than that of Wikilingua. However, they also have a longer MDD than CLCTS hEn. This may be because of the combined effect of (1) shorter sentences from CLCTS hEn (also from Wikilingua En) as discussed above, (2) more conversational texts with shorter dependencies from CLCTS hEn documents (plays and short stories) Wang and Liu (2017), and (3) CLCTS hEn documents mostly from the late 1800s which share more similarities to modern English.
Divergence in lexicon The vocabulary used in historical and modern language varies greatly due to spelling, morphological and semantic changes. For example, in our historical German dataset, the letter “ß” often represents “ss” in modern texts (e.g., “daß” vs. “dass”). Omitting “h”, “t” or “e” is another type of spelling change such as “Thüre” (door, now “Türe” or “Tür”), “thun” (to do, now tun), “Todt” (death, now “Tod”) and “gieng” (passive tense of to go, now “ging”). Such changes in German are a result of two spelling reformations, one in 1901 and one in 1996. In English, we can also observe changes in lexicon such as the word “you” written as “thou” and the morphological change of the second person “will” as “wilt”. To quantify these divergences, we compute the Jaccard similarity (Niwattanakul et al., 2013) of the lexicons between documents and their summaries in Table 4. Historical German has the lowest lexical overlap with modern German, with a Jaccard score of 0.1 for HISTSUMM and 0.19 for CLCTS hDe-De. The pattern is similar for English.
Dataset | Jaccard |
CLCTS hEn-En | 0.234 |
CNN/Daily Mail En-En | 0.377 |
Wikilingua En-En | 0.342 |
CLCTS hDe-De | 0.186 |
HISTSUMM hDe-De | 0.101 |
MLSUM De-De | 0.335 |
Wikilingua De-De | 0.314 |
Semantic shift is another source of language change (Eger and Mehler, 2016; Hamilton, Leskovec, and Jurafsky, 2016). For example, the word “mistress” in Shakespeare’s work has a similar meaning as “sweetheart”. The “car” mentioned in Shakespeare’s work (e.g., “Phoebus’ car”) refers to a special kind of chariot. Another example is the adjective “gay” which is used in the sense of “glad and cheerful”, for example, “The great city in which he dwelt was very gay, for every day strangers visited the town” (The Emperor’s New Clothes by Christian Andersen in 1888). Similarly, the meaning of the word “knave” degenerates from “servant” in the story of Brother Grimm to the modern meaning “a dishonest man” over time.
To conclude: (1) Historical German documents from HISTSUMM and CLCTS contain longer sentences with decreasing sentence length over time (temporal effect) while the sentence length of our historical English documents shows barely any difference to modern texts because of the genre effect. (2) German sentences show an overall higher MDD than that of English sentences in our entire corpus due to the joint effect of genre and publication time. (3) The vocabularies of historical and modern languages differ greatly in terms of spelling, morphological, and semantic changes.
4 Experimental setup
In this section, we introduce the experimental setup of CLCTS. Figure 4 illustrates the model designs (detailed in Section 4.1) and evaluation strategies (detailed in Section 4.2).
4.1 Models
Our experiments consist of three methods: (1) pipeline methods based on extractive summarizers, (2) e2e transformer-based methods with intermediate task finetuning, and (3) zero-shot ChatGPT with different prompt strategies. For finetuned methods (extractive and abstractive), we use 5-fold cross-validation to obtain reliable results for all CLCTS datasets. For ChatGPT, we query summaries for all CLCTS datasets as well to make sure the results are comparable.888During 5-fold cross-validation, we accumulate the test sets from each fold and in this way, we obtain a final test set consisting of all samples. For example, for direction hDe-En with 328 samples, we accumulate output summaries from each test set consisting of 65-66 samples and overall we collect 328 samples after 5-fold cross-validation. In contrast, for zero-shot ChatGPT, we query directly from ChatGPT and collect 328 summaries at once. We summarize model details in Table 5.
Model name | Methods | Type | CV |
MemSum | |||
- | extract-then-translate pipeline with 25 maximal extracted sentences per source document | extractive pipeline | ✓ |
- | extract-normalize-translate pipeline with 25 maximal extracted sentences per source document | ||
mLED | |||
-Base | finetune with CLCTS target task alone | abstractive e2e | ✓ |
-Historical MT | finetune with CLCTS historical translation dataset & CLCTS target task | ||
finetune with CLCTS non-target task & CLCTS target task | |||
mLED | |||
finetune with intermediate tasks listed in Table 6 & CLCTS target task | abstractive e2e | ✓ | |
-MLS-CLS | |||
-MLS-CLS-CTS | |||
ChatGPT as summarizer | |||
-e2e (Title) prompt | prompt in e2e style using only Title and Author | zero-shot abstractive e2e | |
-e2e prompt | prompt in e2e style using source documents | ||
-pipeline prompt | prompt in pipeline style using source documents | ||
- | retrieve max 100 key sentences & prompt in e2e style using retrieved texts | zero-shot abstractive e2e (two-step) |
4.1.1 Pipeline based on extractive summarizer
For extractive summarization, we first use an extract-then-translate pipeline. We choose MemSum (Gu, Ash, and Hahnloser, 2022) as our model, which utilizes multi-step episodic Markov decision processes to extract important sentences from a given text. We use 5-fold cross-validation to train and test MemSum with our monolingual dataset. We limit the maximal extracted sentences per text to 25 according to the average length of summaries in Section 3. During inference, we sort the sentences according to the original text’s order and then translate the extracted sentences to the target language using a multilingual machine translation model developed by Facebook (M2M) Fan et al. (2021). We refer to this model as MemSum-translation. We additionally experiment with historical text normalization before translation. We use the normalization method Norma999https://github.com/comphist/norma Bollmann (2012). We choose the combined version that leverages Lookup, Rule-based, and Distance-based normalization which is claimed to perform well for both English and German Bollmann (2019).101010Norma is recommended by the author when only little training data is available We refer to this model as MemSum-Norma-translation.
4.1.2 Transformer-based e2e models
We modify pretrained models together with intermediate task finetuning to build e2e models. We propose a modified multilingual mBART50 model pretrained for machine translation Tang et al. (2020) with an efficient attention mechanism taken from LED (Beltagy, Peters, and Cohan, 2020) as model architecture. This architecture is suitable for multilingual long sequences as it benefits from the cross-linguality of mBART and the efficiency of processing long documents from LED. We refer to this model as mLED. We use 5-fold cross-validation to obtain reliable results. Due to resource limitations, our mLED models have a maximum of 4,096 input tokens and the number of tokens for the generated summary ranges from a minimum length of 100 tokens to a maximum length of 512 tokens.
Baselines
We first provide baselines of mLED finetuned only within the CLCTS corpus:
for baseline 1, we finetune the model only with the target task; for example, for the model in direction hDe-En, we only finetune mLED with our training set of the same direction. We refer to this model as mLED-Base.
For baseline 2 (baseline 1 + historical machine translation), we utilize our historical translation dataset. We first finetune with the translation task and then with the target task, which we call mLED-Historical MT.
(3) Finally, for baseline 3, we finetune with datasets from both directions. We add prefixes to indicate the directions. We refer to this model as mLED-.
Intermediate task finetuning Apart from the final finetuning step with our target CLCTS task, we conduct intermediate task finetuning using external sources including MLS, CLS, and CTS. Table 6 depicts the detailed design.
Model | MLS | CLS | CTS |
CNN/Daily Mail En-En (tgt) | |||
MLSUM De-De (src) | |||
CNN/Daily Mail En-En (tgt) | |||
MLS-CLS | Wikilingua De-De & En-En | Wikilingua De-En | |
MLS-CLS-CTS | Wikilingua De-De & En-En | Wikilingua De-En | HISTSUMM hDe-En |
(1) Direction hDe-En | |||
Model | MLS | CLS | CTS |
MLSUM De-De (tgt) | |||
Wikilingua En-En (src) | |||
Wikilingua De-De (tgt) | |||
MLS-CLS | Wikilingua En-En & De-De | Wikilingua En-De | |
MLS-CLS-CTS | Wikilingua En-En & De-De | Wikilingua En-De | HISTSUMM hDe-De |
(2) Direction hEn-De |
4.1.3 CLCTS with ChatGPT
We test the summarization ability of ChatGPT with our CLCTS corpus.
Prompt Strategies
We utilize GPT-3.5-turbo
via the OpenAI API to query the summary. We set the temperature to 0.7 to allow for some randomness. We adopt (1) a pipeline translate-then-summarize prompt as well as a (2) e2e prompt. We set prompt lines to the same language as the target output since ChatGPT works better in this setting according to our initial experiments.111111During our first round of data collection all the prompt lines were in English. We noticed that ChatGPT tends to output summaries in the wrong target language, especially when the summary is expected to be in German.
We truncate the texts that exceed ChatGPT input limits.121212According to our experiment with the GPT2 tokenizer, we find that on average our German texts have 1.73 tokens per word (which is higher than the suggested 1.3 tokens per word ratio from OpenAI) and English 1.07 tokens per word. Therefore, we limit the German text to 2,048 words and English to 3,000 words. Since the 4,096 tokens limits from GPT-3.5-turbo are shared between both prompts and completions (outputs), we need to leave some tokens for the output summary.
We also utilize other information such as title and author to test if ChatGPT has explicit “prior” knowledge of the input document and its corresponding summary. We list all our prompt lines in Table 7.
Retrieve-then-summarize with ChatGPT To tackle input limitations of ChatGPT (e.g., for GPT-3.5-turbo a total of 4,096 tokens are shared between prompt and completion),131313https://help.openai.com/en/articles/4936856-what-are-tokens-and-how-to-count-them we build the retrieve-then-summarize design, where we use MemSum as a key sentence retriever and input the retrieved sentences to ChatGPT and request summarization in the target language. For this experiment, we limit the maximal extracted sentences per document to 100 according to the input limits of ChatGPT.
Type | Src-Tgt | Prompt |
e2e | hDe-En | Please summarize the following text in English : [Text]. |
e2e | hEn-De | Bitte fasse den folgenden Text auf Deutsch zusammen: [Text]. |
e2e (Title) | De-En | Please give me the summary of the story [Title] written by [Author]. |
e2e (Title) | En-De | Bitte gebe mir die Zusammenfassung der Geschichte [Titel] von [Autor]. |
pipeline | hDe-En | Please first translate the following text into English and summarize the translated text: [Text] |
pipeline | hEn-De | Bitte übersetze zuerst den folgenden Text auf Deutsch und fasse den übersetzten Text zusammen: [Text]. |
4.2 Evaluation
To identify the best model, we utilize multiple recently popular reference-based metrics. We also conduct a multi-phase human evaluation to evaluate the model and test the effectiveness of the resulting metrics. During the multi-phase evaluation, we give feedback to annotators after each phase for them to better align with the instructions. We list the details of all three evaluation strategies in Table 8.
Type | Aspects | Source | Detail |
Lexical overlap | ROUGE-variant Lin (2004) | ROUGE-1 matching the texts through computing the unigram overlap; ROUGE-L leveraging the longest common subsequence | |
Automatic | Semantic overlap | BERTScore Zhang et al. (2020b) | measuring semantic overlap through calculating the token similarity |
MoverScore Zhao et al. (2019) | measuring semantic overlap utilizing Word Mover’s distance | ||
Generation-based | BARTScore Yuan, Neubig, and Liu (2021) | metric based on text generation using BART | |
NLI-based | MENLI Chen and Eger (2023) | metric based on natural language inference | |
Discourse-based | DiscoScore Zhao, Strube, and Eger (2023) | metric using BERT to model discourse coherence | |
Human | consistency, coherence, fluency, and relevance | SummEval Fabbri et al. (2021) | 6 annotators, 130 annotated instances for direction hDe-En; 104 annotated instances for hEn-De. |
ChatGPT | prompting | - | prompts consisting of the same instruction for human annotators; 494 annotations for hDe-En & 407 for hEn-De. |
Automatic Evaluation
We consider six scores as automatic metrics to reflect five different aspects for measuring the quality of output summaries. The configurations used for each automatic metric are presented in Table 18 (appendix).
Human Evaluation We use six ‘expert’ annotators to evaluate all models. Three of the annotators are female and three are male. All annotators have high competence in English as they are affiliated with the university: they are Master students (3), postgraduates (1), PhD students (1), and NLP faculty members (1). Three of them are native German speakers. Thus, the generated English summaries are evaluated by all six annotators and the generated German summaries are annotated by three native speakers.
We utilize the evaluation strategy by Fabbri et al. (2021) where the annotators score the summary from four perspectives: coherence, consistency, fluency, and relevance. The outputs are rated with a score ranging from 1 (worst) to 5 (best) for each dimension and we allow 0.5 increments. The annotators are presented with the source document, one reference summary, and one output summary during annotation.141414The annotation details including instructions and the number of annotated instances per model are given in Section 9.5 (appendix). After phase 1, we give feedback to the annotators by providing a comparison of their scores with the average scores. The scores that differ substantially are highlighted and we encourage them to review the differences.
To speed up the process, we did not evaluate the same amount of documents for all the models.
Nonetheless, we make sure that each model is evaluated for at least 4 identical source documents. For hDe-En, we collect 130 summary level instances and 104 instances for hEn-De.
ChatGPT evaluation Since human annotation is costly, we follow related work Gao et al. (2023); Chiang and Lee (2023) and test ChatGPT as an evaluator for our CLCTS task. We differ in the following aspects: (1) our CLCTS task is harder to evaluate due to the long length of both reference and generated summary. (2) Our CLCTS task involves multilingual evaluation for both English and German. We give the same instructions to ChatGPT as to the human annotators. For reproducibility purposes, we set the temperature to 0 to reduce the randomness. We collect evaluations for all available generated summaries from documents that have been annotated by humans in the previous step, resulting in 494 ChatGPT annotations for hDe-En and 407 ChatGPT annotations for hEn-De.
5 Results
5.1 Automatic evaluation metrics
The experiment results are shown in Table 9 for hDe-En (upper) and hEn-De (lower).
Model | ROUGE-1/L | BERTScore-P/R/F1 | BARTScore | MoverScore | MENLI-/// | DiscoScore |
Supervised Extractive | ||||||
MemSum | ||||||
- | 0.320/0.175 | 0.521/0.574/0.545 | -3.311 | 0.552 | -0.232/0.407/0.463/0.474 | 0.551 |
- | 0.315/0.171 | 0.514/0.567/0.538 | -3.360 | 0.551 | -0.248/0.398/0.450/0.461 | 0.536 |
Supervised Abstractive - Baseline CLCTS | ||||||
mLED | ||||||
-Base | 0.391/0.201 | 0.537/0.561/0.547 | -3.524 | 0.568 | -0.206/0.418/0.469/0.479 | 1.838 |
-Historical MT | 0.393/0.200 | 0.539/0.566/0.551 | -3.492 | 0.569 | -0.216/0.416/0.473/0.485 | 1.725 |
0.388/0.198 | 0.537/0.561/0.547 | -3.534 | 0.567 | -0.231/0.408/0.465/0.477 | 1.142 | |
Supervised Abstractive - intermediate finetuning | ||||||
mLED | ||||||
0.389/0.198 | 0.540/0.556/0.547 | -3.567 | 0.568 | -0.255/0.398/0.461/0.473 | 1.426 | |
0.386/0.201 | 0.536/0.547/0.54 | -3.593 | 0.568 | -0.252/0.397/0.453/0.464 | 1.595 | |
-MLS-CLS | 0.380/0.198 | 0.530/0.544/0.536 | -3.615 | 0.567 | -0.241/0.400/0.448/0.458 | 1.685 |
-MLS-CLS-CTS | 0.386/0.202 | 0.532/0.548/0.539 | -3.590 | 0.568 | -0.246/0.399/0.451/0.462 | 1.635 |
Zero-shot Abstractive | ||||||
ChatGPT as summarizer | ||||||
-e2e (Title) prompt | 0.304/0.164 | 0.527/0.530/0.528 | -3.857 | 0.542 | -0.700/0.211/0.367/0.398 | 0.788 |
-e2e prompt | 0.399/0.244 | 0.646/0.607/0.624 | -3.363 | 0.575 | -0.194/0.455/0.580/0.605 | 0.982 |
-pipeline prompt | 0.382/0.232 | 0.637/0.597/0.615 | -3.422 | 0.571 | -0.282/0.415/0.554/0.581 | 0.907 |
- | 0.382/0.234 | 0.639/0.597/0.615 | -3.421 | 0.571 | -0.206/0.446/0.566/0.589 | 0.883 |
(a) Direction hDe-En | ||||||
Model | ROUGE-1/L | BERTScore-P/R/F1 | BARTScore | MoverScore | MENLI-/// | DiscoScore |
Supervised Extractive Pipeline | ||||||
MemSum | ||||||
- | 0.315/0.131 | 0.830/0.839/0.834 | -5.320 | 0.814 | -0.648/0.236/0.387/0.418 | 0.301 |
- | 0.311/0.130 | 0.829/0.838/0.833 | -5.350 | 0.814 | -0.652/0.234/0.383/0.413 | 0.291 |
Supervised Abstractive - Baseline CLCTS | ||||||
mLED | ||||||
-Base | 0.321/0.142 | 0.853/0.843/0.848 | -5.275 | 0.823 | -0.605/0.273/0.462/0.499 | 1.653 |
-Historical MT | 0.328/0.143 | 0.853/0.845/0.849 | -5.301 | 0.823 | -0.600/0.276/0.465/0.503 | 1.411 |
0.328/0.144 | 0.854/0.844/0.849 | -5.276 | 0.824 | -0.601/0.276/0.467/0.505 | 1.157 | |
Supervised Abstractive - intermediate finetuning | ||||||
mLED | ||||||
0.316/0.141 | 0.855/0.841/0.848 | -5.253 | 0.826 | -0.613/0.269/0.460/0.498 | 1.500 | |
0.321/0.145 | 0.852/0.841/0.847 | -5.311 | 0.825 | -0.618/0.266/0.453/0.491 | 1.804 | |
-MLS-CLS | 0.328/0.147 | 0.855/0.844/0.850 | -5.250 | 0.826 | -0.586/0.283/0.472/0.510 | 1.050 |
-MLS-CLS-CTS | 0.341/0.153 | 0.855/0.845/0.850 | -5.255 | 0.827 | -0.624/0.268/0.468/0.508 | 0.764 |
Zero-shot Abstractive | ||||||
ChatGPT as summarizer | ||||||
-e2e (Title) prompt | 0.288/0.135 | 0.857/0.840/0.849 | -5.009 | 0.822 | -0.745/0.218/0.444/0.489 | 0.590 |
-e2e prompt | 0.282/0.150 | 0.876/0.843/0.859 | -4.949 | 0.823 | -0.655/0.269/0.509/0.557 | 0.572 |
-pipeline prompt | 0.285/0.147 | 0.874/0.843/0.858 | -4.933 | 0.824 | -0.636/0.274/0.505/0.551 | 0.797 |
- | 0.269/0.143 | 0.874/0.843/0.858 | -4.914 | 0.824 | -0.638/0.274/0.506/0.552 | 0.605 |
(b) Direction hEn-De |
Supervised extractive pipelines
Finding 1. Slightly diminishing results from historical text normalization: We observe a decrease in all metrics scores on MemSum pipeline normalized with Norma for both hDe-En and hEn-De. This is contrary to our hypothesis that spelling normalization of historical texts can boost the translation quality for translators originally trained for modern text. Even though Norma performs well at token-level normalization, it fails to consider the context of the words, where in our use case some historical spellings may correspond to multiple modern meanings depending on the context. An example is the word “in” which can function as an adverb in the phrase for example, “in den Wald” (in the forest) or the accusative case of the word “ihn” (him) in historical texts. A false normalization of such cases results in the meaning change of the entire sentence.
Supervised abstractive e2e Finding 1. Mixed results on intermediate task finetuning: For hDe-En, intermediate task finetuning with external sources does not improve scores, and sometimes finetuning with more tasks even deteriorates the outcome according to the metrics. This may be because our CLCTS hDe-En dataset, which dates back to the early 1800s, has substantial linguistic differences from the external sources as discussed in Section 3. For hEn-De, the results show the opposite. We observe higher scores for models trained with more intermediate summarization tasks from external sources. Though, according to MENLI variants and BARTScore, it is not always better to add CTS tasks. This may be because the CTS tasks from HISTSUMM contain documents that have up to 200 years of difference in publication time and again lead to substantial linguistic differences.
Finding 2. Historical translation task improves model performance only mildly:
For hDe-En, one of the best finetuned models is mLED-Historical MT, where we finetune intermediately with our historical translation task before our target CLCTS task, though according to the metrics, the improvement is very small. Similarly, for the other direction hEn-De, the historical translation task improves model scores but only mildly compared to intermediate tasks. Since the historical translation dataset contains only 201 pairs, a larger size of the historical translation dataset may further improve the results.
Zero-shot abstractive ChatGPT Finding 1. A close match between pipeline and e2e prompts: For hDe-En, we observe slightly better scores for summaries generated with e2e prompts which tell ChatGPT to summarize in the target language directly (see Table 7). However, in the following Section 6, we show that the e2e prompts are prone to invalid outputs.
Finding 2. Competitive results from the retrieve-then-summarize design with ChatGPT: We observe competitive results from retrieve-then-summarize design for both hDe-En and hEn-De. Especially for hDe-En, it is scored as one of the top two models by ROUGE-L, MoverScore, BERTScore and MENLI. Also worth noting is that this design not only performs well but also expands ChatGPT’s potential for long document summarization.
Finding 3. ChatGPT can output summaries from its memory: For the e2e (Title) prompt experiment, we input zero text from the document but only the title, author, and publication year. Slightly surprisingly, for both directions, the resulting scores are mostly comparable to that of the supervised extractive models if not better. ChatGPT is capable of outputting summaries based solely on meta-information for all of our documents. This raises our question on the actual summarization ability of ChatGPT and therefore in the following Section 6,
we design multiple tests to investigate in more detail about ChatGPT involving CLCTS task.
Comparison among all approaches All the metrics give lower scores to extract-then-translate models (with/without normalization) and ChatGPT e2e (Title) prompt than other models, except for BARTScore, which rates the MemSum pipeline as the highest for hDE-En. However, mixed results are observed from the metrics for determining the best models. BERTScore variants in both directions give higher scores to ChatGPT-related models (excluding e2e (Title) prompt). MENLI variants agree with this fully for hDe-En and partially agree for hEn-De. The same is observed from MoverScore and ROUGE-L. BARTScore contradicts with them and scores ChatGPT models higher for hEn-De, though not for hDe-En. Moreover, for hEn-De, ChatGPT is no longer dominating. However, the study from Goyal, Li, and Durrett (2022a) also shows that current metrics fail to evaluate the zero-shot model output properly. To better understand the performance of models and the metrics, we also consider human evaluation.
According to our automatic metrics, we find (1) extractive pipelines perform the worst among all methods; (2) supervised abstractive e2e methods outperform pipeline methods, where the historical translation task improves model performance mildly for both directions and we speculate that the effect of intermediate task finetuning depends on the linguistic similarity of external sources and target sources; (3) zero-shot abstractive ChatGPT delivers the best results under e2e prompts but can output summaries from its memory based on the experiment results from e2e (Title) prompts.
5.2 Human and ChatGPT evaluation
Models | Coh. | Con. | Flu. | Rel. |
hDe-En | 0.497 | 0.595 | 0.150 | 0.572 |
hEn-De | 0.353 | 0.493 | 0.433 | 0.605 |
Table 10 shows the average Spearman’s ranking correlation coefficient among the annotators for coherence, consistency, fluency, and relevance. We obtain moderate agreement for relevance in both directions and moderate agreement for coherence and consistency in hDe-En. The annotation agreement for hEn-De is slightly worse at a weak-moderate level except for relevance. One exception is fluency in hDe-En, where the correlation is poor.
Model | Coherence | Consistency | Fluency | Relevance |
Supervised Extractive | ||||
MemSum | ||||
- | 2.52/1.60 | 2.92/1.50 | 2.98/1.20 | 3.32/1.80 |
- | 2.50/1.50 | 2.92/1.40 | 2.95/1.10 | 3.20/1.70 |
Supervised Abstractive - Baseline CLCTS | ||||
mLED | ||||
-Base | 2.10/1.62 | 2.08/1.12 | 2.78/1.50 | 2.22/1.38 |
-Historical MT | 2.86/1.67 | 2.63/1.56 | 3.01/1.78 | 3.19/1.67 |
- | 2.43/1.30 | 2.28/1.10 | 2.54/1.30 | 2.58/1.20 |
Supervised Abstractive - intermediate finetuning | ||||
mLED | ||||
2.75/1.22 | 2.56/1.11 | 3.08/1.22 | 3.39/1.22 | |
2.38/1.00 | 2.30/1.00 | 2.78/1.10 | 2.75/1.00 | |
-MLS-CLS | 2.20/1.00 | 1.90/1.00 | 2.48/1.00 | 2.42/1.00 |
-MLS-CLS-CTS | 2.11/1.10 | 1.92/1.00 | 2.30/1.10 | 2.28/1.10 |
Zero-shot Abstractive - ChatGPT | ||||
ChatGPT as summarizer | ||||
-e2e (Title) prompt | 2.55/1.70 | 3.05/1.20 | 4.10/3.70 | 2.02/1.40 |
-e2e prompt | 4.14/3.30 | 3.98/2.40 | 4.18/3.80 | 3.97/2.70 |
-pipeline prompt | 4.35/3.30 | 4.30/2.40 | 4.55/4.00 | 4.35/3.00 |
- | 4.08/2.90 | 4.08/2.10 | 4.15/3.50 | 4.25/2.70 |
(a) Direction hDe-En | ||||
Model | Coherence | Consistency | Fluency | Relevance |
Supervised Extractive | ||||
MemSum | ||||
- | 2.47/1.50 | 2.34/1.25 | 2.62/1.38 | 2.19/1.38 |
- | 2.53/1.38 | 2.12/1.25 | 2.75/1.38 | 2.22/1.25 |
Supervised Abstractive - Baseline CLCTS | ||||
mLED | ||||
-Base | 2.01/1.00 | 1.56/1.00 | 2.32/1.00 | 1.65/1.00 |
-Historical MT | 1.98/1.38 | 1.78/1.25 | 2.74/1.38 | 1.78/1.25 |
- | 2.75/1.38 | 2.03/1.12 | 2.78/1.25 | 1.97/1.25 |
Supervised Abstractive - intermediate finetuning | ||||
mLED | ||||
1.99/1.12 | 1.72/1.12 | 2.43/1.12 | 1.91/1.12 | |
1.83/1.00 | 1.70/1.00 | 2.33/1.00 | 1.66/1.00 | |
-MLS-CLS | 1.80/1.00 | 1.71/1.00 | 2.39/1.00 | 1.98/1.00 |
-MLS-CLS-CTS | 2.41/1.00 | 1.88/1.00 | 2.34/1.00 | 2.00/1.00 |
Zero-shot Abstractive - ChatGPT | ||||
ChatGPT as summarizer | ||||
-e2e (Title) prompt | 3.31/2.00 | 2.25/1.00 | 3.28/3.50 | 2.31/1.38 |
-e2e prompt | 3.31/3.00 | 2.80/2.22 | 3.38/3.56 | 2.98/2.33 |
-pipeline prompt | 3.34/2.86 | 3.19/2.00 | 3.47/3.57 | 3.31/2.57 |
- | 3.38/3.00 | 2.94/2.38 | 3.34/3.75 | 3.06/2.62 |
(b) Direction hEn-De | ||||
Models | Coh. | Con. | Flu. | Rel. |
hDe-En | 0.559 | 0.458 | 0.755 | 0.599 |
hEn-De | 0.657 | 0.650 | 0.539 | 0.672 |
(c) Annotation agreement for human and ChatGPT |
We report the mean value of human and ChatGPT (GPT-4-1106-preview) annotations for each model in Table 11.151515We utilize both GPT-3.5-turbo and GPT-4-1106-preview for annotation. GPT-4-1106-preview correlates better with human annotators than GPT-3.5-turbo. We report the annotation results from GPT-4-1106-preview in Table 11 in the main text. The annotation results from GPT-3.5-turbo are reported in Table 21 in Section 9.6 (appendix). The best scores for supervised abstractive models in both directions range between 2 to mildly above 3 by human annotators and below 2 by ChatGPT, which indicates the quality of outputs is bad to mediocre. However, the highest scores for ChatGPT as a summarizer are over 4 (very good) for hDe-En and 3 to 4 (mediocre to good) for hEn-De by human annotators. Overall, the quality of outputs for hEn-De is lower than that for hDe-En. This is also expected since the length of historical English documents is much longer with 9,643 words on average which results in more information loss due to text truncation because of computational limitations161616We also experimented with mBART finetuning with unlimiformer under both low-cost training and computationally expensive long-range training method. Since the low-cost training method only sees full texts at test time and long-range training suffers from the same GPU memory constraints as mLED, we do not observe improvements in model performance compared to baseline mLED. and the cross-lingual embedding distance from Table 3 also suggests more difficulty of hEn-De compared to hDe-En.
For both directions, humans give higher scores to summaries generated by ChatGPT, especially ChatGPT pipeline prompt for hDe-En with an increase of annotation score to a great margin. Unlike the evaluation metrics, humans prefer pipeline prompt output over e2e prompt output for hDe-En and regarding consistency, fluency and relevance for hEn-De. For comparing finetuned models, humans give higher scores for the model finetuned with historical translation in coherence and consistency similar to evaluation metrics for hDe-En. The ratings for hEn-De differ where human ratings prefer mLED-Bidirection. Since the overall quality of generated summaries is worse for hEn-De, the comparisons among outputs may be more difficult for both humans and evaluation metrics.
Mixed results for ChatGPT (GPT-4-1106-preview) as an evaluator. Table 21 (c) shows the annotation agreement between humans and ChatGPT annotation. For both directions, the agreement is moderate to good overall with a slightly better agreement for direction hEn-De (except for fluency). Similarly to the findings from Chiang and Lee (2023), we observe lower Likert scores from ChatGPT evaluation compared to human evaluation except for summaries generated by ChatGPT itself. However, the output for hEn-De is more concerning where we observe more ratings of score 1. Another point worth noticing is that even though in our instruction, we allow for 0.5 increments, we do not observe such ratings from ChatGPT output.
To sum up, ChatGPT as an evaluator (1) is prone to lower scores than humans;171717We include one sample annotation from the gold standard annotations during instruction. (2) cannot distinguish summaries of low quality (hEn-De); (3) does not fully follow the instruction of 0.5 score increments.
5.3 Example outputs
[…]
For low quality generation (summary 1.2 and summary 2.2), we observe a substantial amount of factual inconsistency with the source text (1) and nonsensical text (3). In summary 2.2, we also notice the untranslated expression “Nimmersatt" repeatedly appears in the text. Unclear coreference of pronouns is another issue that affects the readability of the outputs. For example, in the sentence from summary 1.2 “the priest asks him if he (2) can come to the wedding”. The pronoun “he” is ambiguous as it could refer to the boy or the priest himself. In contrast, summary 1.1 and summary 2.1 contain fewer errors, especially the ChatGPT output which only contains one factual inconsistency. In summary 1.1, we find two factual inconsistencies with the source text where the boy “shares” the food instead of “cooking” it himself and he “asks” the image instead of “asking for” it. Additionally, we also notice several other errors such as unclear coreferences of pronouns, nonsensical texts, and spelling mistakes (such as “could’t”).
Reference Texts |
“A poor peasant boy hears a sermon that the way to heaven always leads straight ahead. He goes and comes to the service in a big city church, thinks he is in heaven and cannot be sent away. He sees people praying in front of Mary with the baby Jesus and thinks it is God. He shares his food with him, and the image grows thicker. Once, when he is bedridden for eight days, he goes and apologizes to the image. It invites him to the wedding next Sunday. The priest, who has been watching him, is not allowed to go, but gives him communion on Sunday, when the boy falls down dead.” |
1.1 Generated Text from model mLED-Base, quality level 3-4 |
“A poor peasant boy wants to go to Heaven, but the parson tells him to go straight. When he arrives in the church, he thinks God is there, and says, "Listen, dear God, I will bring you half of my food every day." Every day he cooks (1) half of his food, and the image begins to eat. A couple of weeks later, when people realize that the image is getting bigger and bigger, he stays in church and sees his bread being shared with the mother God. He gets sick and can’t get up for eight days. The parson goes after him and hears him say, "God, don’t take it so long (3), I was sick and could’t stand up." He answers (2), "That doesn’t matter, I’ve seen your good will, that’s enough, and next Sunday you are to come to my wedding." The boy rejoices and asks for the image (1). "No, it’s you alone," says he. He (2) first wants to prepare and gives him the supper. The next Sunday, when supper is coming to him, he falls down and is dead and is to eternal wedding." ” |
1.2 Generated Text from model mLED-MLS-CLS-CTS , quality level 1-2 |
“A poor peasant in the church hears God (1) admonishing him to go to Heaven, and he goes to God’s church. The priest tells him not to go out, but he wants to keep him (1). He stays in church and when he sees the mother God statue, which is cut from wood, the little Jesus (3), he says that it’s God, which means that he will eat half of his food every day (1). After a few weeks, people marvel at the statue and see how he spreads bread with God (1). After eight days, the priest asks him if he (2) can come to the wedding. The boy replies, "No, you’re only me, I told you that you told me that you were you alone. I’ll give him dinner first, which he says he will give him, and then he dies, then he’ll never be able to marry again," so he has to go back to the church and asks if he could come with God alone. After the wedding, the boy asks the parson to give him the dinner, but God says he’s dead and he is dead, and God says that he must marry him again (3).” |
Example 1 |
Reference Texts |
“The wolf is the stronger of the two and dominates the fox. The fox would like to be rid of him. The wolf forces the fox to get him food. He gets him a lamb and leaves. But the wolf is even hungrier and tries to get a lamb for himself. But he is so clumsy that the farmers notice and beat him. The wolf says to the fox that he has tricked him. The fox replied, ‘Why are you such a glutton!’ The next day the fox has to get pancakes for the wolf. They go to the house together and the fox sneaks around the house to carefully get six pancakes from the plate. The wolf again doesn’t have enough and wants to get more. But he again doesn’t manage and the farmer’s wife beats him again. The following day, the two go to the butcher. They jump through a window hatch into the cellar. The wolf eats a lot. The fox, on the other hand, eats little and keeps looking to see if he still fits through the window hatch. Suddenly the butcher comes, the red fox jumps out quickly, only the wolf gets stuck and is beaten to death by the butcher.” |
2.1 Generated Text from model ChatGPT-e2e prompt, quality level 4-5 |
“The wolf and the fox were walking in the woods together. The wolf demanded that the fox bring him food or he would eat him. The fox suggested they steal a lamb from a nearby farm. The wolf ate the lamb and wanted more, but when he tried to steal another, he was caught and beaten by the farmers. The next day, the fox suggested they steal pancakes from a farm, but the wolf ate too many and was caught again. On the third day, the fox suggested they steal salted meat from a cellar, but the wolf ate too much and got stuck in the cellar. The farmers (1) found him and killed him, while the fox escaped and was happy to be rid of the greedy wolf.” |
2.2 Generated Text from model mLED, quality level 1-2 |
“The wolf wants to catch the fox because he is the weakest and would be the master. (1) The fox tells the wolf to get a lamb. (1) The lamb is not happy, (3) but the other wants to have it too. The farmers run to get the lamb, and they hit him (2). When they come to rescue them, the witch tells them they are a nimmersatt (3, 4). They go to a farmer’s house, where they (1) cook pancakes, and the sheep eat the pancakes again. (1) The sheep shout loudly and calls for help, but when she comes out, she says that they have been slaughtered by a man in the forest, and he says that he wants to help out. (3) The wolfe lets go and jumps through the hole, but he does not want to see anything. He jumps out of the hole and kills the shepherd, who does not know what to do. (3) On the third day, he (2) is happy that he is rid of the old Nimmersatt (4), who has been rid of a man who is slaughtering meat in the cellar, and who is the first to get rid of him. (3)” |
Example 2 |
6 Discussion and analysis
6.1 How well do different metrics perform at segment level?
Figure 5 shows Spearman’s ranking correlation between human annotation and evaluation metrics at the segment level. Recall that the correlation results between GPT-4 and human annotation are provided in Table 11 (see Table 21 for GPT-3.5).
Decent correlation from BERTScore variants but worse for German For hDe-En, we observe a weak to moderate level of correlation with human annotation for BERTScore, BARTScore, , and . MoverScore correlates weakly in coherence and relevance and the same is true for ROUGE-L. ROUGE-1 has low correlations with human annotation and the same applies for and DiscoScore. The situation is slightly worse for hEn-De where we observe a lower level of correlation in general compared to hDe-En and less significant correlations from BARTScore and BERTScore variants according to the Student’s t-test especially for coherence. ROUGE-L correlates moderately for direction hEn-De in consistency and relevance. Overall, BERTScore variants correlate slightly better than other metrics with humans in hDe-En for coherence (0.67), consistency (0.61), and relevance (0.66), and in hEn-De for consistency (0.41) and relevance (0.37).
6.2 Statistical analysis on model performance
To quantify the effect of document level characteristics on model performance and contribute to explainability of our models, we analyze how BERTScore-F1 as our best automatic metric overall changes depending on four document features reported in Section 3, namely the mean length of documents (Length), mean dependency distance (MDD), publication year (Year) and divergence of sentence semantics between document and summary by computing the mean sentence-level cosine similarity (Similarity). We apply normalization to numerical features including BERTScore-F1, Length, MDD and Similarity to have zero mean and unit variance.181818We exclude the concern of multicollinearity for numerical variables by computing variance inflation factors (VIF) and the details are reported in Section 9.8 (appendix). We use publication year as a categorical variable where we categorize it by two time periods for hDe-En, namely 1800-1850 (Year:1800-1850) and after 1850 (Year:1850+), and for hEn-De we add another grouping which includes all documents before 1800 (Year:-1800). To examine model effects, we also encode a categorical variable to represent different models (Model). The regression formula is thus:
where represents documents, represents year groups, and represents models defined in Section 4 and
is an error term. We fit Equation (1) using ordinary least squares via statsmodels Seabold and Perktold (2010) separately for both directions (hDe-En and hEn-De).
Direction | Numerical | Categorical | |||
hDe-En | 0.06∗∗∗ | -0.14∗∗∗ | 0.03∗ | base | -0.05 |
hEn-De | 0.24∗∗∗ | -0.15∗∗∗ | -0.08∗∗∗ | -0.003 | 0.69∗∗∗ |
Effect of document features We report the regression results of document features in Table 13.191919The adjusted is 0.425 for direction hDe-En and 0.551 for direction hEn-De. For both directions, embedding similarity measured by has a significantly positive effect on model performance which indicates the model performs better with higher embedding similarity between document and summary. In contrast, the mean length of document (measured by ) shows a significantly negative relation to model performance where more tokens in the source document result in worse performance. This is expected since longer inputs increase the difficulty of summarization and also increase the chance of information loss due to input truncation. The effect of mean dependency distance () varies. We obtain a significantly negative effect for hEn-De which indicates that sentences with more syntactic complexity have a negative influence on model performance. A positive effect for hDe-En is obtained but with a low level of significance. The effect of document publication year is not significant for hDe-En where the model performance differs insignificantly between base group Year:1800-1850 and group Year:1850+. For hEn-De, this also holds for documents published in 1800-1850 compared to the base group with the publication year before 1800. However, we observe a significantly positive effect on model performance if we compare the group Year: 1850+ to base group Year:-1800. This effect also coincides with our finding from Section 3 that English documents from the late 1800s share more similarities to modern English, and thus have a positive impact on model performance.
Also worth noticing is that the main contributor among document features to model performance varies between directions. For hDe-En, the main influence comes from the length of document and mildly from embedding similarity while for hEn-De, the main contributor is the publication year (group Year:1850+) where embedding similarity and the length of document rank the 2nd and 3rd among document features respectively. Overall, however, our findings are largely consistent and they are meaningful: as we argued in the introduction, older texts are more difficult to handle, and the same is true for longer and more syntactically complex source documents; more generally, the more divergence between source documents and reference, the harder is the summarization task.
Effect of model choice
The effect of model choice coincides with our findings from Section 5 where for both directions, we obtain a significant gain with ChatGPT models (except for ChatGPT title prompt) and in contrast, extractive models (extract-then-translate) are significantly worse than the baseline model mLED-Base. For hDe-En, the ChatGPT title prompt model performs significantly worse than the baseline mLED-Base, though not significant for hEn-De. We also notice that for hDe-En, models with intermediate finetuning provide significantly worse performance compared to the baseline as discussed in Section 5 and for hEn-De, the improvement of model performance adding intermediate finetuning tasks is not significant. We include the complete results in Section 9.8 (appendix).
Effect of training set size As discussed in Section 3 and Table 2, the size of our corpora is much smaller compared to other datasets such as CNN/Daily Mail, Wikilingua, and MLSUM. To investigate whether the limited amount of data for finetuning is the main cause of bad performance, we harvest additional instances for both directions. Specifically, we expand the existing datasets for hDe-En from 328 to 455 instances (38.6% increase) and for hEn-De from 289 to 501 instances (73.4% increase). The newly added sources for hDe-En mainly include 95 chapters from the philosophical novel Wilhelm Meister’s Apprenticeship published in 1795, 12 books of epic poetry Reineke Fuchs published in 1794, and 15 scenes from Friedrich Schiller’s play The robbers published in 1781. For hEn-De, the new sources consist of 210 historical translations of fairy tales from different parts of the world published between 1885 and 1910 and 2 short stories.202020The publication time of the newly acquired datasets for both directions closely aligns with that of the original dataset discussed in Section 3. We retrain the CLCTS baseline models using the expanded datasets for both directions, employing the same methods described in Section 4. To ensure comparability, we perform automatic evaluations on the same test sets. Table 19 shows the results for both directions. For hDe-En, the baseline models exhibit marginal improvements, with a mean increase of 0.004 in ROUGE-1, 0.010 in ROUGE-L, and 0.008 in BERTScore-F1. Conversely for hEn-De, the model performances stay invariant on average according to ROUFE-L and BERTScore-F1 with a slight decrease according to ROUGE-1 (-0.004). We speculate that the decrease is because the original CLCTS hEn-De datasets contain more diverse sources (fairy tales, short stories, and plays) while the newly acquired training set primarily contains fairy tales. Moreover, as demonstrated in Section 6.1, the correlation between automatic evaluation metrics and human annotation is worse for hEn-De, which may also affect the evaluation. Overall, our experiments with a larger training set size only show marginal improvements to the existing baseline models. Based on this observation, we hypothesize that a greater volume of data might be necessary to effectively address the intricate nature of the CLCTS task.
6.3 Analysis on ChatGPT summarization
6.3.1 The effect of different prompts and target language
A case study with ChatGPT-e2e vs. ChatGPT-pipeline. Though in Section 5 we observe better metric and human annotation results from ChatGPT e2e prompt, we notice that ChatGPT under this prompt is prone to output summaries in the wrong target language (invalid output). In such cases, we query the summary with the same prompt until a valid output is returned. We calculate the occurrences of invalid outputs after 2 rounds of queries and report them in Table 14, where we observe that using e2e prompts has a higher chance to yield invalid outputs. What is also worth noticing is when the output is expected in German, we observe more invalid cases among all three prompts which means that ChatGPT tends to output English texts when our expected language is in German.
Model | Tgt.lang | Invalid/Obs. |
ChatGPT-title | De | 1/328 |
ChatGPT-pipeline | De | 34/328 |
ChatGPT-e2e | De | 57/328 |
ChatGPT-title | En | 0/289 |
ChatGPT-pipeline | En | 2/289 |
ChatGPT-e2e | En | 8/289 |
6.3.2 ChatGPT for historical text translation
[…]
6.3.3 ChatGPT knows it all?
Recently, the prevalence of LLMs has raised the concern of data contamination Balloccu et al. (2024); Sainz et al. (2023). Since our corpus-building process is highly dependent on Wikipedia (and popular stories presumably discussed on other places on the web), we want to know to what extent ChatGPT has an unfair advantage. We ask ChatGPT to output summaries based on the author, title, and publication year. ChatGPT reports summaries to all requests in both language directions, though the output may contain irrelevant information besides the actual summarization.212121For example, “The story emphasizes the power of kindness and the importance of inner beauty. It also highlights the value of perseverance and the rewards that can come from staying true to oneself.”
However, despite the overall outstanding performance of ChatGPT summarization according to humans, ChatGPT evaluation, and automatic evaluation metrics, ChatGPT as a summarizer can also be prone to hallucinations like other NLG systems Bang et al. (2023); Ji et al. (2023). For example, in the summary of the famous story Cinderella, we notice the golden slippers in the original document become glass slippers which is the best-known version worldwide but is incorrect based on the input text.
Adversarial attacks for ChatGPT Knowing that ChatGPT has specific prior knowledge of our CLCTS task, we go a step further and design three adversarial experiments inspired by the work from Chen and Eger (2023) where we gradually reduce the amount of specific prior knowledge in the input documents (from seen documents to pseudo-unseen documents to unseen documents). We aim to answer (1) does ChatGPT summarize truly based on the input text given its specific prior knowledge? (2) Can ChatGPT summarize against its specific prior knowledge? (3) How well can ChatGPT summarize facing unseen documents?
For question (1), we design a sentence omission task. By dropping an increasing amount of sentences from the original documents, we control the information reduction of ChatGPT input for summarization. We randomly select 11 historical English documents from our CLCTS corpus containing 100-150 sentences and construct new documents by dropping sentences of varying percentages. We then use ChatGPT to summarize the new documents. To quantify the information change, we use the summary of the original documents (generated by ChatGPT) as the reference and evaluate generated summaries from documents after omission using automatic evaluation metrics.222222We set the temperature to 0 since our goal is not to obtain the best summaries but to test ChatGPT summarization performance under gradual information loss. We average the resulting metric scores after scaling and use it as our similarity measure. The result is shown in Figure 6. We observe an overall downward trend of similarity score of both CTS and CLCTS which reflects a certain level of information loss accordingly.
For question (2), we first construct pseudo-unseen documents by entity swap, negation, and omission.232323For each attack, we construct 6 documents. We list examples of adversarial attacks in Section 9.9. For each document, we query the summaries 3 times at temperatures 0, 0.7, and 1 for both English and German. After excluding summaries in the wrong target language, we have 300 valid generated summaries. Entity swap: We swap multiple entities from the original story with new entities. For example, in the story of Cinderella, we switch the hazel tree into a cactus, and (to make it funny) we replace the prince with a watermelon (to ensure coherence, we also add background such as “They live in the Fruit Kingdom.”). Negation: We alter the important aspects of the plot to make a ‘negated story’. For example, in the story of Cinderella, the stepmother and stepsisters are mistreated by the rich man’s evil daughter and the prince eventually marries one of the stepsisters. Omission: For omission, we omit important plots. For example, Cinderella neither attends the ball nor marries the prince. We make sure each pseudo-unseen document only contains one kind of attack. We manually annotate the resulting output where we focus on whether the attack is completed accurately or not (e.g., if during negation, Cinderella is still mistreated by the stepmother — as in the original story — then the task is judged as a failure).242424We obtain the inter-annotator agreement between two annotators of 0.7 according to Cohen’s kappa and 0.87 according to percentage accuracy based on 15 samples. We list examples of failure cases in Section 9.10 (appendix). We compute the accuracy and report it in Table 15.
Attack | CTS | CLCTS |
Omission | 0.79 | 0.67 |
Entity Swap | 0.83 | 0.71 |
Negation | 0.76 | 0.53 |
We observe from the results that (1) ChatGPT can better handle omission and entity swap than negation. This may be because compared to omission and entity swap, negation attacks alter the underlying logical connections of the story and thus require more reasoning to complete this attack under the assumption that ChatGPT is affected by its explicit knowledge during summarization. (2) Higher accuracy for CTS than CLCTS task. For all three attacks, we observe a higher accuracy of completing the task for CTS. This is understandable since CLCTS also involves summarization cross-lingually which is more difficult.
To further investigate question (3), we create 15 unseen documents from fiction chapters whose last updates are in the year 2022 or later252525We include chapters from 5 Chinese novels and 10 English novels.. For modern fiction chapters, we translate and rewrite the stories in historical languages (by adding historical language elements such as morphological changes) and write reference summaries in modern languages. Both documents and summaries are proofread by a co-author of this work. Similar to Section 5, we evaluate the resulting summaries by three human annotators from four perspectives. The results are shown in Table 16. We notice that the overall rating for CTS and CLCTS ranges between 3 and 4 indicating that ChatGPT summarizes unseen documents at a mediocre level which is similar to the results from seen documents. This rating over all four dimensions for hEn-De is slightly worse than the results from our previous experiment using CLCTS corpus (we list the result of the same direction in row 3 in Table 16). We also query and annotate English summaries of unseen documents (CTS hEn-En) where we observe the overall performance is better than CLCTS hEn-De for unseen documents. This is expected since the CTS task does not involve cross-lingual summarization. However, we observe a decrease in summary quality of 0.75 points on average compared to the annotation results from CLCTS hDe-En task from our experiment. This is slightly surprising, since both target languages are in English, and especially CLCTS hDe-En has a more complex task setting. This may again indicate the benefits of the specific prior knowledge from ChatGPT.
Models | Coh. | Con. | Flu. | Rel. |
unseen CLCTS hEn-De | 3.24 | 2.79 | 3.06 | 2.85 |
(our CLCTS hEn-De | 3.31 | 2.80 | 3.38 | 2.98) |
unseen CTS hEn-En | 3.44 | 3.12 | 3.44 | 3.29 |
( our CLCTS hDe-En | 4.14 | 3.98 | 4.18 | 3.97) |
Overall, we argue that ChatGPT can clearly outperform finetuned models for CLCTS tasks in our experiments. However, it does not “magically” solve all the issues, as per our experiments: (1) it profits from its prior knowledge to a certain degree where it may also blend its prior knowledge with the observed data and ChatGPT is better at omission and entity swap than negating against its prior knowledge; (2) the benefit from prior knowledge slightly inflates its assessed quality where ChatGPT generates summaries of mediocre quality from unseen source documents with a slightly worse performance over all four evaluating dimensions compared to seen documents for hEn-De; (3) the performance is language dependent where compared to German outputs (hEn-De), English outputs (hEn-En) yield slightly better ratings.
7 Concluding remarks
In this work, we build the first CLCTS corpus and a historical translation dataset with additional information for future explorations. We study the characteristics of this corpus both alone and in comparison to other summarization corpora and examine the effectiveness of pipeline models, popular e2e transformer-based abstractive models leveraging intermediate task finetuning and ChatGPT for the CLCTS task. Even though these methods have been shown effective for single-task settings such as CLS or long document summarization, they fail to deliver good summaries in our combined task setting (cross-lingual, cross-temporal, and long document). We showcase the potential of GPT-3.5 in CLCTS as a zero-shot summarizer. It provides moderate to good quality outputs and seems very adept as a context-aware spelling normalization tool. However, it profits from its prior knowledge to a certain degree as tested by adversarially attacked and unseen source documents. Overall, we observe better performance for plot omission and entity swap than plot negation against its prior knowledge. Moreover, GPT-3.5 performs slightly worse for unseen source documents compared to seen documents.
As for evaluation, ChatGPT (both GPT-3.5 and GPT-4) as an evaluator can correlate with human evaluations at a moderate to good level but is prone to giving lower scores. Additionally, we find a moderate correlation of BERTScore with human annotation (slightly worse for German).
By regression analysis with BERTScore-F1, we quantify the effects of document features on model performance where we find positive impacts from documents with more recent publication years and document-summary embedding similarity while we observe a negative relation between model performance and the length of document. This helps explain model performance and affirms our hypotheses about the difficulty of the CLCTS task as discussed in the introduction (language change and divergence, etc.).
Future work can address the limitations revealed by our experiments. For example, researchers can explore other recent models of great potential, such as Unlimiformer under sufficient GPU memory Bertsch et al. (2024) and Longnet Ding et al. (2023). Other training frameworks such as pretraining under multi-task learning are also promising alternatives. Incorporating tasks such as non-parallel historical corpora for pretraining and the PAR3 cross-temporal dataset for machine translation Thai et al. (2022) may further provide benefits. Finally, future work can focus on extending the size and diversity of our CLCTS datasets, including more diverse language pairs.262626Our current work builds the CLCTS corpus for the high-resource language pair German and English. Main difficulties working with other languages include: (1) Even more Wikipedia summaries may be missing for other languages (German and English are two of the highest resource languages); (2) it may be more difficult to obtain historical sources of other languages; (3) human evaluation of other languages may be less accessible and thus even more costly; (4) automatic metrics including GPT4 may also be worse as we have observed worse performance for German than for English. While exploring more distant language pairs is thus immensely interesting, (5) our results further indicate that transfer would be even more difficult for all models involved.
8 Acknowledgments
We thank all annotators for their hard work and time. The NLLG group gratefully acknowledges support from the Federal Ministry of Education and Research (BMBF) via the grant “Metrics4NLG” and the German Research Foundation (DFG) via the Heisenberg grant EG375/5-1.